### v1

# GaussianFormer3D: Multi-Modal Gaussian-based Semantic Occupancy Prediction with 3D Deformable Attention

Abstract
3D semantic occupancy prediction is critical for achieving safe and reliable autonomous driving. Compared to camera-only perception systems, multi-modal pipelines, especially LiDAR-camera fusion methods, can produce more accurate and detailed predictions. Although most existing works utilize a dense grid-based representation, in which the entire 3D space is uniformly divided into discrete voxels, the emergence of 3D Gaussians provides a compact and continuous object-centric representation. In this work, we propose a multi-modal Gaussian-based semantic occupancy prediction framework utilizing 3D deformable attention, named as GaussianFormer3D. We introduce a voxel-to-Gaussian initialization strategy to provide 3D Gaussians with geometry priors from LiDAR data, and design a LiDAR-guided 3D deformable attention mechanism for refining 3D Gaussians with LiDAR-camera fusion features in a lifted 3D space. We conducted extensive experiments on both on-road and off-road datasets, demonstrating that our GaussianFormer3D achieves high prediction accuracy that is comparable to state-of-the-art multi-modal fusion-based methods with reduced memory consumption and improved efficiency. Project website: GaussianFormer3D.
footnotetext: Corresponding author.
1 Introduction
Perception systems are essential for constructing safe and reliable autonomous vehicles [19, 89]. Among various perception tasks, 3D semantic occupancy prediction [5, 24, 22, 88, 70, 36] is particularly crucial as it enables fine-grained understanding of both semantics and geometry information of the environments. Recent advances in vision-based occupancy prediction have demonstrated impressive performance on large-scale datasets [4, 1, 39, 62]. However, camera sensors are sensitive to lighting conditions and lack accurate depth estimation, motivating researchers to incorporate other sensor modalities to enhance the robustness of autonomous driving perception systems.
Figure 1: We propose a LiDAR-camera fusion-based semantic occupancy prediction framework named GaussianFormer3D. We use 3D Gaussians instead of dense grids to reduce memory consumption and enhance algorithm efficiency. GaussianFormer3D achieves comparable performance to state-of-the-art multi-modal occupancy methods with reduced memory usage.
LiDAR sensors have been widely applied to autonomous driving for perception tasks such as 3D object detection [93, 77, 81, 59, 8, 30]. Compared to cameras, LiDAR provides more accurate depth information and captures finer geometric relationships of objects, making LiDAR particularly advantageous for 3D semantic occupancy prediction [57, 58, 76, 10, 80, 73, 95, 46, 6, 75, 71]. However, LiDAR-based pipelines may struggle to capture accurate semantic information for small objects, where camera-based methods excel. To balance semantics and geometry, multi-modal fusion-based algorithms have been proposed to leverage the strengths of different sensors. Fusion-based semantic occupancy prediction methods include LiDAR-camera fusion [52, 66, 32, 3, 63, 60, 87], camera-radar fusion [72, 45] and LiDAR-camera-radar fusion [51]. Among these sensor configurations, LiDAR-camera fusion is the most popular and top-performing one.
Most existing LiDAR-camera occupancy networks employ a 3D voxel-based [52, 66, 32, 87] or a 2D bird’s-eye-view (BEV)-based representation [60], both depicting a 3D scene as a dense grid-based structure. Despite achieving comparable performance, they inevitably struggle with redundant empty grids and high computational costs. Recently, inspired by the success of 3D Gaussian splatting [28], a novel object-centric representation has been explored for the first time in vision-based 3D semantic occupancy prediction [25, 23, 13]. GaussianFormer [25, 23] represents a 3D scene as a set of 3D Gaussians, each consisting of a mean, covariance and semantic label. These Gaussians are refined with a 2D deformable attention mechanism [94], and then processed by an efficient Gaussian-to-voxel splatting module to predict the semantic occupancy. However, current Gaussian-based methods [25, 23] rely solely on 2D image feedback to update 3D Gaussians, limiting their ability to model 3D space with accurate depth information and fine-grained geometric structure. How to exploit data from other modalities, such as LiDAR, to refine and obtain a more accurate 3D Gaussian representation for efficient semantic occupancy prediction remains unexplored.
Based on the aforementioned observations, we propose GaussianFormer3D: a multi-modal Gaussian-based semantic occupancy prediction framework with 3D deformable attention, as shown in Fig. 1. GaussianFormer3D models a scene using 3D Gaussians initialized from LiDAR voxel features, updates Gaussians through 3D deformable attention in a LiDAR-camera unified 3D feature space, and finally predicts semantic occupancy via Gaussian-to-voxel splatting. To the best of our knowledge, GaussianFormer3D is the first multi-modal semantic occupancy network that employs a Gaussian-based object-centric scene representation. In summary, our main contributions are as follows:
  • We propose a novel multi-modal Gaussian-based semantic occupancy prediction framework. By integrating LiDAR and camera data, our method significantly outperforms camera-only baselines with similar memory usage.
  • We design a voxel-to-Gaussian initialization module to provide 3D Gaussians with geometry priors from LiDAR data. We also develop an enhanced 3D deformable attention mechanism [31] to update Gaussians by aggregating LiDAR-camera fusion features in a lifted 3D space.
  • We present extensive evaluations on two on-road datasets, nuScenes-SurroundOcc [70] and nuScenes-Occ3D [64], and one off-road dataset, RELLIS3D-WildOcc [83]. Results show that our method performs on par with state-of-the-art dense grid-based methods while having reduced memory consumption and improved efficiency.
2 Related Work
Figure 2: An overview of the proposed GaussianFormer3D framework. We first voxelize LiDAR point clouds to obtain non-empty voxel features for initializing the position and opacity of 3D Gaussians [28]. Then multi-scale LiDAR depth maps and camera feature maps are extracted through projection and an image backbone respectively, and multiplied via outer product to construct a lifted 3D fusion feature space. Gaussians are iteratively updated with 3D sparse convolution, 3D deformable attention, and property refinement. The Gaussian representation is eventually processed by a Gaussian-to-voxel splatting module [25] to generate dense 3D semantic occupancy.
Multi-Modal Semantic Occupancy Prediction. Multi-modal occupancy methods generally outperform single-modal ones since different modalities can complement each other. Common multi-modal sensor configurations for semantic occupancy include LiDAR-camera [52, 66, 32, 3, 63, 60, 87], camera-radar [72, 45] and LiDAR-camera-radar [51]. Among them, LiDAR-camera is the top-performing one as it combines LiDAR’s accurate depth and geometry sensing with camera’s powerful semantic perception. Similar to single-modal pipelines, existing LiDAR-camera occupancy networks mainly build on voxel-based [52, 66, 32, 87, 68] or BEV-based representations [60]. CONet [68] proposes a coarse-to-fine pipeline to sample 3D voxel features for refining the coarse occupancy prediction. Co-Occ [52] obtains multi-modal voxel features through a geometric and semantic-aware fusion module, and employs a NeRF-based implicit volume rendering regularization [49] to enhance the fused representation. OccGen [66] and OccMamba [32] encode multi-modal inputs to produce voxel fusion features, and then decode the features using diffusion denoising [17] and hierarchical Mamba modules [14] respectively to predict semantic occupancy. OccFusion [87] transforms LiDAR and camera inputs into multi-modal voxel features via 2D deformable attention [94] followed by an occupancy head.
3D Gaussians for Autonomous Driving. Due to the inherent advantages of modeling scenes explicitly and continuously, 3D Gaussians [28] have been adopted as the scene representation over the traditional grid-based solutions in 3D semantic occupancy prediction [25, 23, 13] and 4D semantic occupancy forecasting [96]. 3D Gaussians also demonstrated their superiority in real-time image rendering and novel view synthesis and thus have been adopted for driving scene reconstruction and simulation [92, 56, 33, 78, 16]. Furthermore, end-to-end autonomous driving [90] and visual pre-training [74, 85] utilize 3D Gaussians as the driving world representation for various downstream perception and planning tasks. However, these approaches are mainly designed for camera-only autonomous driving, neglecting the potential of multi-modal data in Gaussian initialization and updating. GSPR [55] proposes a Gaussian-based multi-modal place recognition algorithm, and SplatAD [16] designs the first 3D Gaussian splatting pipeline to render both LiDAR and camera data. In this work, we explore utilizing multi-modal data, especially from LiDAR and camera, to learn a better 3D Gaussian representation for more accurate and efficient semantic occupancy prediction.
3 Method
The overview of GaussianFormer3D is presented in Fig. 2.
3.1 Scene as 3D Gaussian Representation
Semantic occupancy prediction aims to understand both semantic information and geometric structure of the environment. In the multi-modal scenario, given multi-view camera images =$\mathcal{I}=\{\mathbf{I}_{i}\}_{i=1}^{N_{c}}$ and LiDAR point cloud $\mathcal{P}=\{\mathbf{P}_{i}\}_{i=1}^{N_{p}}$, $\mathbf{P}_{i}=(x_{i},y_{i},z_{i},\eta_{i})$ containing the 3D position and intensity of the point, the goal is to predict the semantic occupancy grid $\mathbf{O}\in\mathcal{C}^{X\times Y\times Z}$, where $N_{c}$, $N_{p}$, $\mathcal{C}$ and $X\times Y\times Z$ represent the number of camera views, the number of LiDAR points, the set of semantic classes and the size of the voxel grid, respectively. Unlike uniform grids in traditional dense grid-based representations, the 3D Gaussian representation can adaptively depict the regions of interest due to the universal approximation capability of Gaussian mixtures [25, 11]. Specifically, a scene is modeled as a set of 3D Gaussians $\mathcal{G}=\{\mathbf{G}_{i}\}_{i=1}^{N_{g}}$, where each Gaussian $\mathbf{G}_{i}$ is parameterized by its mean $\mathbf{m}_{i}\in\mathbb{R}^{3}$, rotation $\mathbf{r}_{i}\in\mathbb{R}^{4}$, scale $\mathbf{s}_{i}\in\mathbb{R}^{3}$, opacity $\sigma_{i}\in[0,1]$ and semantic label $\mathbf{c}_{i}\in\mathbb{R}^{|\mathcal{C}|}$. $N_{g}$ is the total number of Gaussians representing the scene. The value of Gaussian $\mathbf{G}$ evaluated at location $\mathbf{x}$ can be calculated as:
$\mathbf{g}(\mathbf{x};\mathbf{G})=\sigma\cdot{\rm{exp}}\big{(}-\frac{1}{2}(\mathbf{x}-\mathbf{m})^{\rm T}\mathbf{\Sigma}^{-1}(\mathbf{x}-\mathbf{m})\big{)}\mathbf{c},$ (1)
$\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\rm T}\mathbf{R}^{\rm T},\quad\mathbf{S}={\rm{diag}}(\mathbf{s}),\quad\mathbf{R}={\rm{q2r}}(\mathbf{r}),$ (2)
[…]
$\hat{\mathbf{o}}(\mathbf{x};\mathcal{G})=\sum_{i=1}^{N_{g}}\mathbf{g}_{i}(\mathbf{x};\mathbf{m}_{i},\mathbf{s}_{i},\mathbf{r}_{i},\sigma_{i},\mathbf{c}_{i}).$ (3)
Gaussian-to-voxel splatting module is designed to only aggregate Gaussians within a neighborhood of the targeted voxel instead of querying all Gaussians in a scene to improve efficiency and reduce unnecessary computation and storage [25]. Thus, Eq. 3 can be furthur approximated by replacing $N_{g}$ with $N_{g}(\mathbf{x})$, where $N_{g}(\mathbf{x})$ is the number of neighboring Gaussians at location $\mathbf{x}$. During training, the Gaussian-based occupancy model is trained in an end-to-end manner, supervised by the ground truth semantic occupancy label $\mathbf{\bar{O}}\in\mathcal{C}^{X\times Y\times Z}$. Both cross entropy loss $L_{ce}$ and the lovasz-softmax loss $L_{lov}$ [2] are used for supervision.
3.2 Voxel-to-Gaussian Initialization
Two sets of 3D Gaussian features are adopted in our model following GaussianFormer [25]. The first set consists of learnable Gaussian physical properties $\mathcal{G}=\{\mathbf{G}_{i}\in\mathbb{R}^{d}\}_{i=1}^{N_{g}}$ introduced in Sec. 3.1, where $d=11+|\mathcal{C}|$, which are also our learning targets. The second set is non-learnable high-dimensional Gaussian features $\mathcal{Q}=\{\mathbf{Q}_{i}\in\mathbb{R}^{m}\}_{i=1}^{N_{g}}$, where $m$ is the feature dimension, serving as queries for the attention mechanism [65, 7, 94] and implicitly encode the spatial and semantic information during the Gaussian update. Previous work [25] randomly initializes the Gaussian physical properties during training, and optimizes these properties iteratively through multiple repetitive refinement modules. This design constrains Gaussians to learn complex 3D geometry information solely from 2D images, which will inevitably encounter inaccurate spatial modeling.
To resolve this issue, we propose a LiDAR-based voxel-to-Gaussian initialization strategy to initialize the mean and opacity of Gaussians with geometry priors from LiDAR data, as indicated in the dashed blue box in Fig. 2. Specifically, we first aggregate the most recent $N_{f}$ LiDAR scans into a combined point cloud $\mathcal{\bar{P}}=\{\mathcal{P}_{i}\}_{i=1}^{N_{f}}$. We voxelize the combined point cloud and compute the feature of each non-empty voxel as the mean position and intensity of all points within it [93]. These LiDAR-based voxel features are then used to initialize the position and opacity of 3D Gaussians:
$\mathbf{m}_{i}=\frac{1}{|\mathcal{P}_{v}|}\sum_{j\in\mathcal{P}_{v}}(x_{j},y_{j},z_{j}),\quad\sigma_{i}=\frac{1}{|\mathcal{P}_{v}|}\sum_{j\in\mathcal{P}_{v}}\eta_{j},$ (4)
where $i\in\{1,...,N_{g}\}$ denotes the index of Gaussians to be initialized and $v\in\{1,...,N_{v}\}$ denotes the index of all non-empty voxels; $\mathcal{P}_{v}$ is the set of LiDAR points in $\mathcal{\bar{P}}$ falling into voxel $v$. When $N_{g}<N_{v}$, we randomly choose a subset of non-empty voxels with size of $N_{g}$ to initialize Gaussians, otherwise, we randomly select a subset of Gaussians with size of $N_{v}$ to be initialized with non-empty voxels. After initialization, we apply a 3D sparse convolution module to 3D Gaussians for self-encoding. The information and interactions of these Gaussians are efficiently extracted and aggregated through the sparse convolution operation for updating the Gaussian queries.
3.3 LiDAR-Guided 3D Deformable Attention
Lift, Splat, Shoot (LSS) [54] and 2D attention-based methods [94, 7] are widely adopted for feature lifting which transform multi-view 2D images into a 3D space to obtain lifted features. However, LSS suffers from excessive computational costs, hindering its application to multi-scale feature maps that are important for recognizing objects of various sizes. GaussianFormer [25] utilizes a 2D deformable attention to extract visual information from 2D images. Despite its efficiency, the 2D deformable attention-based method struggles with depth ambiguity. As multiple 3D reference points of different Gaussians can be projected to the same 2D position with similar sampling points in the 2D view, leading to the aggregation of the same 2D features for different 3D Gaussian queries. The underlying reason for this is the lack of accurate depth information during the feature lifting and aggregating process.
A 3D deformable attention operator, namely DFA3D [31], is designed to mitigate the depth ambiguity problem by first expanding 2D feature maps into 3D using estimated depth [12] and then applying an attention mechanism [65] to aggregate features from the expanded 3D feature maps. However, the operator was originally designed for BEV-based 3D object detection, and relies on DepthNet [12] to estimate monocular depth. Inspired by DFA3D [31], we propose a LiDAR-guided 3D deformable attention mechanism for Gaussian-based semantic occupancy prediction, as illustrated in the dashed orange box in Fig. 2. We first form a unified LiDAR-camera 3D feature space $\mathbf{F}^{{\rm 3D}}$ by conducting outer product between the multi-scale depth maps $\mathbf{F}^{\rm d}$, generated from the LiDAR point cloud, and the multi-scale camera feature maps $\mathbf{F}^{\rm c}:\mathbf{F}^{\rm 3D}=\mathbf{F}^{\rm d}\otimes\mathbf{F}^{\rm c}$. For feature sampling, we design a two-stage key point sampling method to aggregate sufficient informative features for updating Gaussian queries. First, we sample a group of 3D reference points $\mathcal{R}_{G}=\{\mathbf{m}_{i}=\mathbf{m}+\Delta\mathbf{m}_{i}|i=1,...,N_{R_{1}}\}$ for each Gaussian $\mathbf{G}$ by shifting its mean $\mathbf{m}$ with learned offsets $\Delta\mathbf{m}$. Then we project these 3D reference points into the fusion feature space $\mathbf{F}^{\rm 3D}$ with extrinsics $\mathcal{T}$ and intrinsics $\mathcal{K}$, where each projected reference point is positioned at $\mathbf{\bar{m}}_{i}=(u_{i},v_{i},d_{i})$. After projection, we further generate learnable sampling offsets $\Delta\mathbf{\bar{m}}_{ij}=(\Delta u_{ij},\Delta v_{ij},\Delta d_{ij})$ for each projected reference point $\mathbf{\bar{m}}_{i}$. The overall sampling points of a given Gaussian $\mathbf{G}$ in the fusion feature space $\mathbf{F}^{\rm 3D}$ can be formulated as:
[…]
where $N_{R_{1}}$ and $N_{R_{2}}$ denote the number of sampling reference points for each Gaussian and for each projected 3D reference point, respectively. Finally, we update the Gaussian query $\mathbf{Q}$ with the weighted sum of aggregated LiDAR-camera fusion features $\Delta\mathbf{Q}$:
$\Delta\mathbf{Q}=\frac{1}{N_{c}}\sum_{c=1}^{N_{c}}\sum_{i=1}^{N_{R_{1}}}\sum_{j=1}^{N_{R_{2}}}{\rm DFA}(\mathbf{Q},\boldsymbol{\pi}_{c}(\mathbf{\bar{m}}_{ij};\mathcal{T},\mathcal{K}),\mathbf{F}_{c}^{\rm 3D}),$ (6)
where $\rm DFA(\cdot)$ and $\boldsymbol{\pi}_{c}(\cdot)$ represent the 3D deformable attention operation and the transformation from the Gaussian coordinate frame to $\mathbf{F}_{c}^{{\rm 3D}}$ coordinate frame generated from camera view $c$ respectively. After acquiring sufficient 3D geometric and semantic information through sparse convolution and 3D deformable attention, the Gaussian query $\mathbf{Q}$ is passed to a multi-layer perceptron (MLP), and decoded to refine the Gaussian property $\mathbf{G}$. We iteratively optimize the Gaussian properties with $B$ blocks of sparse convolution, 3D deformable attention, and refinement modules.
Method Modality IoU $\uparrow$ mIoU $\uparrow$
barrier
bicycle
bus
car
const. veh.
motorcycle
pedestrian
traffic cone
trailer
truck
drive. surf.
other flat
sidewalk
terrain
manmade
vegetation
MonoScene [5] C 24.0 7.3 4.0 0.4 8.0 8.0 2.9 0.3 1.2 0.7 4.0 4.4 27.7 5.2 15.1 11.3 9.0 14.9
BEVFormer [37] C 30.5 16.8 14.2 6.6 23.5 28.3 8.7 10.8 6.6 4.1 11.2 17.8 37.3 18.0 22.9 22.2 13.8 22.2
TPVFormer [22] C 30.9 17.1 16.0 5.3 23.9 27.3 9.8 8.7 7.1 5.2 11.0 19.2 38.9 21.3 24.3 23.2 11.7 20.8
OccFormer [88] C 31.4 19.0 18.7 10.4 23.9 30.3 10.3 14.2 13.6 10.1 12.5 20.8 38.8 19.8 24.2 22.2 13.5 21.4
SurroundOcc [70] C 31.5 20.3 20.6 11.7 28.1 30.9 10.7 15.1 14.1 12.1 14.4 22.3 37.3 23.7 24.5 22.8 14.9 21.9
C-CONet [68] C 26.1 18.4 18.6 10.0 26.4 27.4 8.6 15.7 13.3 9.7 10.9 20.2 33.0 20.7 21.4 21.8 14.7 21.3
FB-Occ [38] C 31.5 19.6 20.6 11.3 26.9 29.8 10.4 13.6 13.7 11.4 11.5 20.6 38.2 21.5 24.6 22.7 14.8 21.6
GaussianFormer [25] C 29.8 19.1 19.5 11.3 26.1 29.8 10.5 13.8 12.6 8.7 12.7 21.6 39.6 23.3 24.5 23.0 9.6 19.1
GaussianFormer-2 [23] C 31.7 20.8 21.4 13.4 28.5 30.8 10.9 15.8 13.6 10.5 14.0 22.9 40.6 24.4 26.1 24.3 13.8 22.0
LMSCNet [58] L 36.6 14.9 13.1 4.5 14.7 22.1 12.6 4.2 7.2 7.1 12.2 11.5 26.3 14.3 21.1 15.2 18.5 34.2
L-CONet [68] L 39.4 17.7 19.2 4.0 15.1 26.9 6.2 3.8 6.8 6.0 14.1 13.1 39.7 19.1 24.0 23.9 25.1 35.7
M-CONet [68] L+C 39.2 24.7 24.8 13.0 31.6 34.8 14.6 18.0 20.0 14.7 20.0 26.6 39.2 22.8 26.1 26.0 26.0 37.1
Co-Occ [52] L+C 41.1 27.1 28.1 16.1 34.0 37.2 17.0 21.6 20.8 15.9 21.9 28.7 42.3 25.4 29.1 28.6 28.2 38.0
Ours L+C 43.3 27.1 26.9 15.8 32.7 36.1 18.6 21.7 24.1 13.0 21.3 29.0 40.6 23.7 27.3 28.2 32.6 42.3
Table 1: 3D semantic occupancy prediction results on SurroundOcc [70] validation set. The best is bolded and the second best is underlined. Our method achieves comparable overall performance with the state-of-the-art LiDAR-camera fusion-based methods, while surpasses them on classes of small objects, dynamic vehicles, and large surfaces.
Method Modality mIoU $\uparrow$
others
barrier
bicycle
bus
car
const. veh.
motorcycle
pedestrian
traffic cone
trailer
truck
drive. surf.
other flat
sidewalk
terrain
manmade
vegetation
MonoScene [5] C 6.1 1.8 7.2 4.3 4.9 9.4 5.7 4.0 3.0 5.9 4.5 7.2 14.9 6.3 7.9 7.4 1.0 7.7
BEVDet [21] C 11.7 2.1 15.3 0.0 4.2 13.0 1.4 0.0 0.4 0.13 6.6 6.7 52.7 19.0 26.5 21.8 14.5 15.3
BEVFormer [37] C 23.7 5.0 38.8 10.0 34.4 41.1 13.2 16.5 18.2 17.8 18.7 27.7 49.0 27.7 29.1 25.4 15.4 14.5
TPVFormer [22] C 28.3 6.7 39.2 14.2 41.5 47.0 19.2 22.6 17.9 14.5 30.2 35.5 56.2 33.7 35.7 31.6 20.0 16.1
CTF-Occ [64] C 28.5 8.1 39.3 20.6 38.3 42.2 16.9 24.5 22.7 21.1 23.0 31.1 53.3 33.8 38.0 33.2 20.8 18.0
RenderOcc [53] C 26.1 4.8 31.7 10.7 27.7 26.5 13.9 18.2 17.7 17.8 21.2 23.3 63.2 36.4 46.2 44.3 19.6 20.7
GaussianFormer* [25] C 35.5 8.8 40.9 23.3 42.9 49.7 19.2 24.8 24.4 22.5 29.4 35.3 79.0 36.9 46.6 48.2 38.8 33.1
BEVDet4D* (2f) [20] C 39.3 9.3 47.1 19.2 41.5 52.2 27.2 21.2 23.3 21.6 35.8 38.9 82.5 40.4 53.8 57.7 49.9 45.8
COTR* (2f) [44] C 44.5 13.3 52.1 32.0 46.0 55.6 32.6 32.8 30.4 34.1 37.7 41.8 84.5 46.2 57.6 60.7 52.0 46.3
PanoOcc* (4f) [69] C 42.1 11.7 50.5 29.6 49.4 55.5 23.3 33.3 30.6 31.0 34.4 42.6 83.3 44.2 54.4 56.0 45.9 40.4
FB-Occ* (16f) [38] C 42.1 14.3 49.7 30.0 46.6 51.5 29.3 29.1 29.4 30.5 35.0 39.4 83.1 47.2 55.6 59.9 44.9 39.6
OccFusion* [87] L+C 48.7 12.4 51.8 33.0 54.6 57.7 34.0 43.0 48.4 35.5 41.2 48.6 83.0 44.7 57.1 60.0 62.5 61.3
Ours* L+C 46.4 9.8 50.0 31.3 54.0 59.4 28.1 36.2 46.2 26.7 40.2 49.7 79.1 37.3 49.0 55.0 69.1 67.6
Table 2: 3D semantic occupancy prediction results on Occ3D [64] validation set. * denotes training with camera visibility mask. (xf) denotes the number of history image frames used for temporal fusion. The best is bolded and the second best is underlined.
4 Experiments
4.1 Datasets
[…]
RELLIS-3D [27] dataset is a multi-modal off-road driving dataset containing RGB images, LiDAR point clouds, stereo images, GPS and IMU data. WildOcc [83] provides the first off-road occupancy annotation on the RELLIS-3D, which are split into 7399/1249/1399 frames for training, validation and testing respectively. The annotation is in the range of $[-20\rm{m},0\rm{m}]\times[-10\rm{m},10\rm{m}]\times[-2\rm{m},6\rm{m}]$, where each voxel has a resolution of $0.2\mathrm{m}$ and labeled as one of 9 classes (7 semantics, 1 other class and 1 empty). WildOcc [83] is used to evaluate the performance of our model on off-road scenes with a LiDAR-monocular setting.
4.2 Implementation and Evaluation Details
For camera branch, we set the resolution of input images as $900\times 1600$ for nuScenes [4] and $1200\times 1920$ for RELLIS-3D [27]. We utilize the ResNet101-DCN [15] checkpoint pretrained from FCOS3D [67] as the backbone and FPN [40] as the neck. For LiDAR branch, we aggregate and voxelize previous 10 sweeps of point clouds, and obtain the mean features through a voxel feature encoder [93]. The LiDAR depth map is generated and saved before training following [31, 35]. The number of Gaussians is set to 25,600 in our main experiments. We employ these Gaussians to only model the occupied space, and leave the empty space to one fixed large Gaussian to improve efficiency [23]. We train our model with an AdamW [42] optimizer with a weight decay of 0.01. The learning rates are set to $1\times 10^{-4}$ for nuScenes and $3\times 10^{-4}$ for RELLIS-3D, and decay with a cosine annealing schedule. Our model is trained for 24 epochs with a batch size of 8 on nuScenes and 20 epochs with a batch size of 4 on RELLIS-3D on Nvidia A40 GPUs. We use Intersection-over-Union (IoU) and mean Intersection-over-Union (mIoU) for evaluation metrics following MonoScene [5]. See supplementary material for the calculation details of the metrics.
Method Mod. IoU $\uparrow$ mIoU $\uparrow$
Grass
Tree
Bush
Puddle
Mud
Barrie
Rubble
Test Set Class Percentage % 41.052 36.094 17.621 0.512 0.774 0.001 0.001
C-OFFOcc (4f)  [83] C 29.7 11.2 24.6 23.8 22.1 0.6 3.5 0.6 3.2
GaussianFormer [25] C 19.5 6.3 21.8 12.1 5.2 2.7 2.3 0.0 0.0
M-OFFOcc [83] L+C - 12.9 - - - - - - -
M-OFFOcc (4f) [83] L+C 32.8 14.8 28.6 33.4 27.5 0.9 6.8 1.7 4.6
Ours L+C 33.9 13.1 24.0 45.4 12.9 6.6 2.8 0.0 0.0
Validation Set Class Percentage % 31.739 42.210 18.497 0.105 0.842 2.218 3.836
GaussianFormer [25] C 23.0 8.2 19.4 24.4 5.2 0.0 4.4 0.0 4.0
Ours L+C 29.5 13.1 19.1 38.5 10.6 0.1 4.6 4.2 14.5
Table 3: 3D semantic occupancy prediction results on WildOcc [83] validation and test sets. We show the percentages of classes in the ground truth occupied voxels for validation and test sets respectively. Results except for GaussianFormer [25] and ours are reported in WildOcc [83]. (xf) denotes the number of history image frames used for temporal fusion. The best is bolded.
Method Mod. Query Form Query Number Lat. (ms) $\downarrow$ Mem. (GB) $\downarrow$ IoU $\uparrow$ mIoU $\uparrow$
BEVFormer [37] C 2D BEV 200$\times$200 310 4.5 30.5 16.8
TPVFormer [22] C 3D TPV 200$\times$(200+16+16) 320 5.1 30.9 17.1
SurroundOcc [70] C 3D Voxel 200$\times$200$\times$16 340 5.9 31.5 20.3
25600 227 4.7 28.7 16.0
GaussianFormer [25] C 3D Gaussian 144000 370 6.1 29.8 19.1
6400 313 3.0 30.4 19.9
12800 323 3.0 30.4 19.9
GaussianFormer-2 [23] C 3D Gaussian 25600 357 3.0 31.0 20.3
50$\times$50$\times$4 532 7.6 33.3 21.2
M-CONet [68] L+C 3D Voxel 100$\times$100$\times$8 670 7.8 39.2 24.7
Co-Occ [52] L+C 3D Voxel 100$\times$100$\times$8 580 11.8 41.1 27.1
6400 415 4.9 39.6 21.4
12800 462 5.0 41.4 24.2
Ours L+C 3D Gaussian 25600 555 5.5 43.3 27.1
Table 4: Efficiency comparison of different methods and effect of Gaussian number on SurroundOcc [70] validation set. The results of our method are tested on one A40 GPU with batch size one during inference. The best is bolded within each modality.
4.3 Quantitative Results
3D semantic occupancy prediction performance. We report the performance of GaussianFormer3D on SurroundOcc [70], Occ3D [64] and WildOcc [83] in Tab. 1, Tab. 2 and Tab. 3, respectively. For on-road scenarios in Tab. 1 and Tab. 2, our method surpasses GaussianFormer [25] extensively on all classes, leading to overall 13.5 and 8.0 increases on the IoU and mIoU respectively on SurroundOcc [70] and 10.9 increase on the mIoU on Occ3D [64]. Compared to state-of-the-art LiDAR-camera approaches [52, 68, 87], GaussianFormer3D achieves comparable overall performance while showing superior performance in predicting small objects (e.g., motorcycle, pedestrian), dynamic vehicles (e.g., car, construction vehicle, truck) and surrounding surfaces (e.g., manmade, vegetation) which are crucial classes for autonomous driving tasks. This improvement is due to Gaussians’ universal approximating ability to model objects with flexible scales and shapes. For off-road results in Tab. 3, our method with single-frame image input surpasses M-OFFOcc [83] using 4 sequential images by 1.1 in IoU and performs on par in mIoU. Moreover, our method outperforms GaussianFormer [25] by 14.4 in IoU and 6.8 in mIoU on the test set, highlighting LiDAR’s role in understanding the geometry of complex off-road terrains. GaussianFormer3D excels in predicting regions with large surfaces, such as grass, tree, and puddle, while remaining suboptimal for subtle terrain variations like mud. For barrier and rubble, their low occurrence in the test set (0.001% of occupied voxels) poses a challenge due to the lack of sufficient features for reliable prediction. See supplementary material for more analysis.
Evaluation of model efficiency and effect of Gaussian number. We evaluate and compare the latency and memory consumption of our approach with other methods in Tab. 4. Our model achieves multi-modal fusion-based prediction performance while maintaining approximately the same low memory usage as camera-only methods. Compared to Co-Occ [52], our method saves about 50% memory consumption, making it more suitable for running onboard on autonomous vehicles. Additionally, our approach employs only 25,600 Gaussians with 28 channels while Co-Occ [52] requires 80,000 queries with 128 channels to achieve similar performance demonstrating the potential of our method to enable more efficient communication for connected vehicles or multi-robot collaborations. The latency of our method is higher than that of vision-based pipelines, which is mainly due to the computation overhead introduced by 3D deformable attention operations. We also examine the effect of the number of Gaussians on the model performance in Tab. 4. As the number of Gaussians increases, both latency and memory consumption rise, while the IoU and mIoU metrics are steadily improved.
4.4 Ablation Study
To break down the performance improvement brought by two designed modules, we conduct extensive ablation experiments to validate our design choices. The main ablation study is conducted in Tab. 5. We observe that both the proposed voxel-to-Gaussian initialization and the LiDAR-guided 3D deformable attention modules contribute to the superior performance of GaussianFormer3D. The voxel-to-Gaussian initialization significantly improves the model’s ability to detect both small objects (e.g., pedestrian, traffic cone) and large surfaces (e.g., manmade structures, vegetation). This validates the effectiveness of multi-sweep LiDAR scans in providing Gaussians with accurate geometric information of occupied space. We also notice that LiDAR-guided 3D deformable attention mechanism enhances the model’s prediction ability on dynamic vehicles (e.g., bicycle, bus, car, motorcycle, trailer, truck) and near-road surfaces (e.g., drivable surface, flatten area, sidewalk, terrain) where objects detected by LiDAR points are visible to surrounding cameras. In these regions, the LiDAR points and corresponding image pixels are associated in the lifted 3D feature space, enabling the model to retrieve aggregated fusion features of on-road and near-road objects.
Figure 3: Qualitative results on the on-road SurroundOcc [70] validation set. Our multi-modal Gaussian-based occupancy method can capture both semantics information and geometry structure of the surroundings. Best viewed on screen and in color.
Model V2G DFA IoU $\uparrow$ mIoU $\uparrow$
barrier
bicycle
bus
car
const. veh.
motorcycle
pedestrian
traffic cone
trailer
truck
drive. surf.
other flat
sidewalk
terrain
manmade
vegetation
29.2 18.8 18.8 11.6 24.6 29.4 10.2 14.8 12.3 8.6 11.9 21.1 39.5 23.6 24.3 22.4 9.5 18.8
40.7 25.8 25.3 17.1 30.9 35.0 17.9 21.5 23.9 14.8 20.3 27.7 37.8 20.4 24.9 25.3 30.1 39.4
40.7 26.4 25.3 17.4 32.4 35.7 17.8 23.9 22.1 12.0 20.5 29.1 41.8 24.6 28.1 27.7 27.5 36.6
GaussianFormer3D 43.3 27.1 26.9 15.8 32.7 36.1 18.6 21.7 24.1 13.0 21.3 29.0 40.6 23.7 27.3 28.2 32.6 42.3
Table 5: Ablation study of proposed modules evaluated on SurroundOcc [70] validation set. Voxel-to-Gaussian and LiDAR-Guided 3D Deformable Attention are abbreviated as V2G and DFA respectively.
[…]
36.7 22.4
34.9 21.2
V2G 40.7 25.8
(a) Ablation study of initialization strategy for V2G. PM-Point denotes probabilistic modeling with point cloud reported in [23].
[…]
40.1 25.0
40.6 25.2
V2G 40.7 25.8
(b) Ablation study of LiDAR voxel size for V2G. The unit of length is $\mathrm{m}$. We set the height of all the voxels as $0.2\mathrm{m}$.
[…]
37.7 24.5
40.1 26.1
DFA 40.7 26.4
(c) Ablation study of offset sampling methods for DFA. We run experiments with applying learnable offset sampling before and after projecting Gaussians into the lifted 3D feature space.
[…]
36.1 22.2
36.6 22.1
DFA 40.7 26.4
(d) Ablation study of feature lifting and aggregating methods for DFA. We concatenate LiDAR sparse and dense depth maps with RGB features respectively to conduct 2D deformable attention-based Gaussian update.
Table 6: Ablation study of module design choices on the SurroundOcc [70] validation set. Voxel-to-Gaussian and LiDAR-Guided 3D Deformable Attention are abbreviated as V2G and DFA respectively.
Figure 4: Qualitative results on the off-road WildOcc [83] test set. Our multi-modal Gaussian-based occupancy method can outperform the ground truth (as shown in the first row) and predict classes such as puddle that are vital for off-road autonomous driving (as shown in the second row). Best viewed on screen and in color.
Figure 5: Visualization comparison with GaussianFormer [25] on SurroundOcc [70]. By incorporating LiDAR, our method can obtain Gaussians with more adaptive scales and shapes, resulting in more accurate semantic predictions and delicate geometry details.
Figure 6: Multi-resolution occupancy prediction using same 3D Gaussians. Leveraging the continuity of Gaussians, our method enables multi-resolution prediction, yielding more accurate and smoother occupancy at higher resolution in certain regions.
Voxel-to-Gaussian Initialization. We first compare different levels of LiDAR features used for initializing Gaussian properties in Tab. 6(a). The improvement achieved with the multi-sweep voxel feature is significantly greater than that of the single-sweep point feature and the point cloud probabilistic modeling strategy used in GaussianFormer-2 [23], which validates the effectiveness of our proposed module. We further conduct an ablation study on the size of LiDAR voxel in initialization in Tab. 6(b). As the voxel size decreases, the model performance slightly improves. We choose $0.075\rm{m}\times 0.075\rm{m}\times 0.2\rm{m}$ as the final size.
LiDAR-Guided 3D Deformable Attention. We first study the effect of the two-stage offset sampling strategy in Tab. 6(c). We observe that applying learnable offset sampling both before and after projection achieves higher performance than single-stage sampling, which validates that our two-stage sampling method can aggregate sufficient informative features for refining Gaussians. We also compare different feature aggregating methods for deformable attention in Tab. 6(d), including 3D deformable attention, 2D deformable attention with concatenated LiDAR sparse depth map and with completed dense depth map [29]. The results validate our design choice of 3D deformable attention.
4.5 Qualitative Results
We visualize 3D Gaussians and occupancy to qualitatively verify the effectiveness of GaussianFormer3D for on-road scenes in Fig. 3. Our method can accurately predict both semantics and fine-grained geometry of the surrounding environments. In some cases, it even outperforms the ground truth by correctly completing occupancy in regions that lack semantic annotations. Qualitative results of our method on off-road scenes are given in Fig. 4. Our method is able to predict semantic occupancy for classes like mud and puddle, which are essential for achieving safe and effective off-road autonomous driving. We further compare our approach with GaussianFormer [25] in Fig. 5. The Gaussians in our method are more adaptive in scales and shapes, precisely appearing in the occupied regions of objects in both long-range and short-range areas, aided by the LiDAR sensor. Additionally, compared to voxel-based discretized approaches that train and predict at a fixed resolution, our method can predict multi-resolution semantic occupancy without additional training cost, attributed to the continuous property of Gaussians. This property enables more accurate and smoother prediction for certain areas when inferred at a higher resolution, as demonstrated in Fig. 6. Please see more qualitative results in the supplementary material.
5 Conclusion
In this paper, we proposed GaussianFormer3D, a novel multi-modal semantic occupancy prediction framework that builds on 3D Gaussian scene representation and 3D deformable attention. We introduced a voxel-to-Gaussian initialization strategy to endow 3D Gaussians with accurate geometry priors from LiDAR data. We also designed a LiDAR-guided 3D deformable attention mechanism to refine 3D Gaussians with LiDAR-camera fusion feature in a lifted 3D space. Extensive experiments show the effectiveness of GaussianFormer3D in achieving accurate and fine-grained semantic occupancy prediction. In the future, we plan to explore the application of our multi-modal 3D Gaussian scene representation for multi-robot coordination.
Supplementary Material
Appendix A Related Work
Single-Modal Semantic Occupancy Prediction. Semantic occupancy prediction, also known as semantic scene completion, was firstly proposed to predict the occupancy and semantic states of all voxels within a pre-defined range for indoor scenarios [86, 34, 61, 9]. The vision-based and LiDAR-based pipelines are first explored, and further extended to outdoor environments, especially the autonomous driving scenes [4, 1, 39, 62]. The key to this challenging task is to learn an informative and compact scene representation, to which the common solutions mainly include 3D voxel-based representation [93, 77], 2D BEV-based representation [37, 79] and TPV-based (Tri-Perspective View) representation [22]. Voxel-based methods utilize discretized voxels to represent a 3D space, and transform semantic and geometric information from 2D images [5, 26, 36, 43, 48, 69, 70, 88, 18, 91, 47, 84, 53] or 3D point clouds [58, 76, 57, 10, 80, 73, 46, 6, 41, 75, 50] to voxel feature vectors. These approaches can effectively capture complicated structures but incur high storage and computational costs due to the dense nature of voxel-based representations. Moreover, a significant portion of these costs is unnecessary given the abundance of empty voxels in autonomous driving scenes. BEV-based representation has also been applied to semantic occupancy prediction [38, 82, 24] after its success in 3D object detection [37, 79]. TPV-based representation [22] extends BEV with two additional perpendicular planes and has been used in both vision-based [24, 22] and LiDAR-based [95] semantic occupancy networks. Although BEV-based and TPV-based methods mitigate the redundancy issue caused by the empty grids, their compression schemes inevitably lead to information loss. 3D Gaussian splatting [28] provides an expressive and compact scene representation, which has been explored in vision-based 3D semantic occupancy prediction [25, 23, 13]. GaussianFormer [25] represents an autonomous driving scene as a set of 3D Gaussians, designed to only model the object-occupied regions, which significantly improved prediction efficiency and reduced memory consumption.
Appendix B WildOcc Dataset
Figure 7: Distribution of categories across the WildOcc [83] train, validation, and test sets. The bar charts illustrate the percentage of each class within the respective dataset splits.
Figure 8: Qualitative results of GaussianFormer3D on the WildOcc [83] validation and test sets. In the off-road scenes, our model can still capture the geometry structures and semantic information of large continuous surfaces (rubble, grass) or small irregular objects (barrier, mud), as shown in the red boxes. Best viewed on screen and in color.
[…]
Based on the split files from RELLIS-3D [27], we filter out samples in the WildOcc that lack corresponding semantic occupancy labels, LiDAR poses, or ground truth depth, obtaining 7399, 1249 and 1399 samples for training, validation and testing, respectively. The test set originally contains 11 classes in total. However, to ensure comparability with the baselines in [83], we follow the same class selection strategy, considering semantic classes including grass, tree, bush, barrier, puddle, mud, and rubble, while grouping the remaining labels into the others category. The percentage distribution of each class is presented in Fig. 7.
A key challenge of this dataset is the significant distribution shift between the training/validation sets and the test set. Notably, as shown in Fig. 7, the barrier and rubble classes exhibit extremely low occurrence in the test set compared to other categories. Specifically, the test set contains only 273 and 233 voxels for barrier and rubble, respectively, accounting for merely 0.0008% and 0.0007% of the occupied voxels. This severe class imbalance poses challenges for model performance, as limited training data can hinder the accurate prediction of these underrepresented categories. Additionally, the insufficiency of captured features further complicates the prediction process, making it even more challenging to reliably infer these classes.
Appendix C Evaluation Metrics
For evaluation metrics, we utilize Intersection-over-Union (IoU) and mean Intersection-over-Union (mIoU) following MonoScene [5]. The IoU and mIoU can be computed as:
$\mathrm{IoU}$ $=\frac{TP_{\neq c_{0}}}{TP_{\neq c_{0}}+FP_{\neq c_{0}}+FN_{\neq c_{0}}},$ (7)
$\mathrm{mIoU}$ $=\frac{1}{|\mathcal{C}^{\prime}|}\sum_{i\in\mathcal{C}^{\prime}}{\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}},$ (8)
where $TP$, $FP$, $FN$, $c_{0}$ and $\mathcal{C}^{\prime}$ represent the number of true positive, false positive, false negative predictions, the empty class and the set of non-empty classes respectively.
Figure 9: Qualitative results of GaussianFormer3D on the Occ3D [64] validation set. Our method is able to recover the fine-grained geometry structures and accurate semantics of the on-road scenes, and even outperforms the ground truth by completing the occupancy at some areas without annotations. Best viewed on screen and in color.
Figure 10: Multi-resolution occupancy predicted by the same 3D Gaussians on the SurroundOcc [70] validation set. The regions within the red boxes show that the high-resolution prediction generated by the same 3D Gaussians can achieve more accurate and smoother results, compared to the low-resolution one and the occupancy ground truth. Best viewed on screen and in color.
Appendix D Implementation Details
For multi-view images, we only apply photometric distortion augmentation, image normalization and image padding. During evaluation, we do not apply any test time augmentation technique. For the 3D deformable attention, DFA3D [31] has proved that the trilinear interpolation in the 3D feature space can be transformed into a depth-weighted bilinear interpolation, which means that in implementation, 3D deformable attention can be transformed into a depth-weighted 2D deformable attention to simultaneously maintain theoretical equivalence and improve efficiency. Hence we take advantage of the depth-weighted 2D deformable attention operator implemented in CUDA by DFA3D [31] to conduct 3D deformable attention.
Appendix E Supplementary Experiments
E.1 Qualitative Results
Qualitative results on the WildOcc [83] validation and test sets. We present visualization results of our model on the WildOcc [83] validation and test sets in Fig. 8. Two representative samples from both the validation and test sets are selected for illustration. Notably, all sequences in the validation and test sets are completely unseen during training. The class distribution has been reported in Appendix B. In the first validation sample (top row), our method successfully predicts the rubble region (highlighted in red boxes), which corresponds to the pile of rocks in the camera view. In the second validation sample, our model effectively captures the barrier structures in the scene. These results indicate that our model can learn semantic occupancy even with limited training data. In the test set samples (bottom two rows), our method accurately reconstructs large continuous surfaces, such as grass and trees, and effectively detects irregular objects and terrains, such as bush and mud, as shown in the red boxes. Overall, our predictions exhibit a high degree of similarity to the ground truth, especially in preserving the geometric structures of various terrain types, including grass, tree, rubble, and barrier.
Qualitative results on Occ3D [64] validation set. We provide visualization results of our model on the Occ3D [64] validation set in Fig. 9. Our method achieves fine-grained semantic occupancy prediction, and even occasionally outperforms the ground truth. Even though the occupancy ground truth is sparse and missing at some regions, our model is able to accurately recovering semantic information of these unlabeled areas.
Figure 11: Visualization comparison of different Gaussian initialization strategies on the SurroundOcc [70] validation set. With LiDAR voxel features as priors, our 3D Gaussians precisely appear at the regions of interests, which paves the way for the following 3D deformable attention-based Gaussian update.
[…]
Method Modality Sunny Rainy Day Night Sunny Rainy Day Night
GaussianFormer [25] C 29.6 27.5 30.3 19.5 18.9 18.0 19.2 9.3
GaussianFormer3D L+C 43.6 (+14.0) 41.6 (+14.1) 43.6 (+13.3) 40.5 (+21.0) 27.3 (+8.4) 25.2 (+7.2) 27.4 (+8.2) 15.5 (+6.2)
Table 7: 3D semantic occupancy prediction results on SurroundOcc [70] validation set for different weather and lighting conditions. We quantitatively show the improvement made by GaussianFormer3D over the baseline GaussianFormer.
Multi-resolution semantic occupancy prediction on SurroundOcc validation set. We provide more qualitative results of multi-resolution semantic occupancy prediction in Fig. 10. As shown in red boxes in Fig. 10, we observe that the high-resolution prediction achieves more accurate and smoother occupancy at some regions. Although Gaussians are supervised with the fixed-resolution occupancy ground truth, multi-resolution predictions can still be generated by the same Gaussian representation due to its continuous modeling property. This advantage eradicates the need of training several models to predict occupancy of different resolutions, and keeps the high prediction accuracy with the same Gaussians, which significantly shortens the algorithm deployment time and saves the computation resources.
Visualization comparison of different Gaussian initialization strategies. We provide qualitative results of visualization comparison of different Gaussian initialization strategies in Fig. 11. Compared to random initialization, our designed voxel-to-Gaussian (V2G) initialization strategy is able to endow the 3D Gaussians with accurate position and geometry priors from the LiDAR data at the beginning of training, which paves the way for the following 3D deformable attention-based Gaussian refinement.
E.2 Quantitative Results
Model performance under different weather conditions. To break down the performance improvement under different weather conditions, we group the scenes in the nuScenes [4] dataset based on climate and lighting conditions. We demonstrate the results of both GaussianFormer3D and GaussianFormer [25] in Tab. 7. Compared to the baseline GaussianFormer, our approach presents increased performance across all weather conditions, which validates the effectiveness and robustness of our model. Due to the introduction of LiDAR sensor, GaussianFormer3D shows a significant performance improvement over the camera-only baseline under extreme climate (rainy) and low lighting condition (night). Both experiments are conducted with 25,600 Gaussians for a fair comparison.

### v2

@@
 Abstract
-3D semantic occupancy prediction is critical for achieving safe and reliable autonomous driving.
-Compared to camera-only perception systems, multi-modal pipelines, especially LiDAR-camera fusion methods, can produce more accurate and detailed predictions.
-Although most existing works utilize a dense grid-based representation, in which the entire 3D space is uniformly divided into discrete voxels, the emergence of 3D Gaussians provides a compact and continuous object-centric representation.
-In this work, we propose a multi-modal Gaussian-based semantic occupancy prediction framework utilizing 3D deformable attention, named as GaussianFormer3D.
-We introduce a voxel-to-Gaussian initialization strategy to provide 3D Gaussians with geometry priors from LiDAR data, and design a LiDAR-guided 3D deformable attention mechanism for refining 3D Gaussians with LiDAR-camera fusion features in a lifted 3D space.
-We conducted extensive experiments on both on-road and off-road datasets, demonstrating that our GaussianFormer3D achieves high prediction accuracy that is comparable to state-of-the-art multi-modal fusion-based methods with reduced memory consumption and improved efficiency.
-Project website: GaussianFormer3D.
-footnotetext: Corresponding author.
-1 Introduction
-Perception systems are essential for constructing safe and reliable autonomous vehicles [19, 89].
-Among various perception tasks, 3D semantic occupancy prediction [5, 24, 22, 88, 70, 36] is particularly crucial as it enables fine-grained understanding of both semantics and geometry information of the environments.
-Recent advances in vision-based occupancy prediction have demonstrated impressive performance on large-scale datasets [4, 1, 39, 62].
-However, camera sensors are sensitive to lighting conditions and lack accurate depth estimation, motivating researchers to incorporate other sensor modalities to enhance the robustness of autonomous driving perception systems.
-Figure 1: We propose a LiDAR-camera fusion-based semantic occupancy prediction framework named GaussianFormer3D.
-We use 3D Gaussians instead of dense grids to reduce memory consumption and enhance algorithm efficiency.
-GaussianFormer3D achieves comparable performance to state-of-the-art multi-modal occupancy methods with reduced memory usage.
-LiDAR sensors have been widely applied to autonomous driving for perception tasks such as 3D object detection [93, 77, 81, 59, 8, 30].
-Compared to cameras, LiDAR provides more accurate depth information and captures finer geometric relationships of objects, making LiDAR particularly advantageous for 3D semantic occupancy prediction [57, 58, 76, 10, 80, 73, 95, 46, 6, 75, 71].
-However, LiDAR-based pipelines may struggle to capture accurate semantic information for small objects, where camera-based methods excel.
-To balance semantics and geometry, multi-modal fusion-based algorithms have been proposed to leverage the strengths of different sensors.
-Fusion-based semantic occupancy prediction methods include LiDAR-camera fusion [52, 66, 32, 3, 63, 60, 87], camera-radar fusion [72, 45] and LiDAR-camera-radar fusion [51].
-Among these sensor configurations, LiDAR-camera fusion is the most popular and top-performing one.
-Most existing LiDAR-camera occupancy networks employ a 3D voxel-based [52, 66, 32, 87] or a 2D bird’s-eye-view (BEV)-based representation [60], both depicting a 3D scene as a dense grid-based structure.
-Despite achieving comparable performance, they inevitably struggle with redundant empty grids and high computational costs.
-Recently, inspired by the success of 3D Gaussian splatting [28], a novel object-centric representation has been explored for the first time in vision-based 3D semantic occupancy prediction [25, 23, 13].
-GaussianFormer [25, 23] represents a 3D scene as a set of 3D Gaussians, each consisting of a mean, covariance and semantic label.
-These Gaussians are refined with a 2D deformable attention mechanism [94], and then processed by an efficient Gaussian-to-voxel splatting module to predict the semantic occupancy.
-However, current Gaussian-based methods [25, 23] rely solely on 2D image feedback to update 3D Gaussians, limiting their ability to model 3D space with accurate depth information and fine-grained geometric structure.
-How to exploit data from other modalities, such as LiDAR, to refine and obtain a more accurate 3D Gaussian representation for efficient semantic occupancy prediction remains unexplored.
-Based on the aforementioned observations, we propose GaussianFormer3D: a multi-modal Gaussian-based semantic occupancy prediction framework with 3D deformable attention, as shown in Fig.
+3D semantic occupancy prediction is essential for achieving safe, reliable autonomous driving and robotic navigation.
+Compared to camera-only perception systems, multi-modal pipelines, especially LiDAR-camera fusion methods, can produce more accurate and fine-grained predictions.
+Although voxel-based scene representations are widely used for semantic occupancy prediction, 3D Gaussians have emerged as a continuous and significantly more compact alternative.
+In this work, we propose a multi-modal Gaussian-based semantic occupancy prediction framework utilizing 3D deformable attention, namely GaussianFormer3D.
+We introduce a voxel-to-Gaussian initialization strategy that provides 3D Gaussians with accurate geometry priors from LiDAR data, and design a LiDAR-guided 3D deformable attention mechanism to refine these Gaussians using LiDAR-camera fusion features in a lifted 3D space.
+Extensive experiments on real-world on-road and off-road autonomous driving datasets demonstrate that GaussianFormer3D achieves state-of-the-art prediction performance with reduced memory consumption and improved efficiency.
+Project website: https://lunarlab-gatech.github.io/GaussianFormer3D/.
+I Introduction
+Perception systems are essential for the development of safe, reliable and intelligent autonomous vehicles and field robots [46].
+Among various perception tasks, 3D semantic occupancy prediction is particularly crucial as it enables fine-grained understanding of both geometric and semantic information of the environments [31].
+For autonomous driving (AD), recent advances in vision-based occupancy prediction have shown impressive results on large-scale datasets [2, 1].
+However, the sensitivity of cameras to lighting variations and their limited depth accuracy still underscore the need to incorporate additional sensing modalities for robust AD.
+LiDAR sensors have been widely applied to AD for perception tasks such as 3D object detection [51, 15].
+Compared to cameras, LiDAR provides more accurate depth information and finer geometric relationships of objects, making it particularly advantageous for 3D semantic occupancy prediction [30, 5, 40, 39, 43, 24].
+However, LiDAR-based pipelines often struggle to capture accurate semantics for small objects, where camera-based methods excel [49].
+To balance geometric accuracy and semantic richness, multi-modal fusion algorithms have been proposed to leverage the strengths of complementary sensing modalities.
+Current approaches include mainly LiDAR-camera fusion [26, 34, 16, 45] and camera–radar fusion [23], with LiDAR-camera fusion showing superior performance and becoming the dominant choice.
+Most LiDAR-camera occupancy networks employ a 3D voxel-based [26, 34, 16, 45, 4] representation to model a scene as a dense grid-based structure.
+Despite achieving comparable performance, these methods inevitably suffer from redundant empty grids and high computational costs.
+Recently, inspired by the success of 3D Gaussian splatting [13], an object-centric Gaussian representation has been explored in vision-based 3D semantic occupancy prediction, achieving improved computational efficiency.
+GaussianFormer series [11, 9] represent a scene as a set of 3D Gaussians, each consisting of a mean, covariance and semantics.
+These Gaussians are refined using a 2D deformable attention [52], and then processed by an efficient Gaussian-to-voxel splatting module to predict semantic occupancy.
+Despite high efficiency, current Gaussian-based methods [11, 9] rely solely on 2D image to update 3D Gaussians, limiting their ability to model 3D space with accurate depth and fine-grained geometric structure.
+How to effectively leverage other sensor modalities, such as LiDAR, to refine and obtain a more accurate 3D Gaussian representation for efficient semantic occupancy prediction remains unexplored and challenging.
+Figure 1: We propose a new LiDAR-camera fusion-based semantic occupancy prediction framework using 3D Gaussians.
+We evaluate it on both on-road and off-road driving scenarios.
+Our method demonstrates superior performance on overall occupancy Intersection-of-Union (IoU), achieves substantial performance gains on small objects (pedestrian, motorcycle) and large surfaces (manmade, vegetation), and consumes less memory during inference.
+In this work, we propose GaussianFormer3D: a multi-modal Gaussian-based semantic occupancy prediction framework with 3D deformable attention, as shown in Fig.
 1.
-GaussianFormer3D models a scene using 3D Gaussians initialized from LiDAR voxel features, updates Gaussians through 3D deformable attention in a LiDAR-camera unified 3D feature space, and finally predicts semantic occupancy via Gaussian-to-voxel splatting.
-To the best of our knowledge, GaussianFormer3D is the first multi-modal semantic occupancy network that employs a Gaussian-based object-centric scene representation.
+GaussianFormer3D models a scene using 3D Gaussians initialized from LiDAR voxel features, updates Gaussians through 3D deformable attention in a LiDAR-camera unified 3D feature space, and predicts semantic occupancy via Gaussian-to-voxel splatting.
+To the best of our knowledge, our model is the first multi-modal semantic occupancy network that employs an object-centric Gaussian-based scene representation.
 In summary, our main contributions are as follows:
   • We propose a novel multi-modal Gaussian-based semantic occupancy prediction framework.
-By integrating LiDAR and camera data, our method significantly outperforms camera-only baselines with similar memory usage.
-  • We design a voxel-to-Gaussian initialization module to provide 3D Gaussians with geometry priors from LiDAR data.
-We also develop an enhanced 3D deformable attention mechanism [31] to update Gaussians by aggregating LiDAR-camera fusion features in a lifted 3D space.
-  • We present extensive evaluations on two on-road datasets, nuScenes-SurroundOcc [70] and nuScenes-Occ3D [64], and one off-road dataset, RELLIS3D-WildOcc [83].
-Results show that our method performs on par with state-of-the-art dense grid-based methods while having reduced memory consumption and improved efficiency.
-2 Related Work
-Figure 2: An overview of the proposed GaussianFormer3D framework.
-We first voxelize LiDAR point clouds to obtain non-empty voxel features for initializing the position and opacity of 3D Gaussians [28].
-Then multi-scale LiDAR depth maps and camera feature maps are extracted through projection and an image backbone respectively, and multiplied via outer product to construct a lifted 3D fusion feature space.
-Gaussians are iteratively updated with 3D sparse convolution, 3D deformable attention, and property refinement.
-The Gaussian representation is eventually processed by a Gaussian-to-voxel splatting module [25] to generate dense 3D semantic occupancy.
+By integrating LiDAR and camera data, ours significantly outperforms camera-only baselines with similar memory usage.
+  • We design a voxel-to-Gaussian initialization module to provide 3D Gaussians with geometry priors from LiDAR, and also develop an enhanced 3D deformable attention mechanism to update Gaussians by aggregating LiDAR-camera fusion features in a lifted 3D space.
+  • We conduct extensive evaluations on two on-road datasets, nuScenes-SurroundOcc [38] and nuScenes-Occ3D [32], along with an off-road dataset, RELLIS3D-WildOcc [44].
+Results show that ours outperforms state-of-the-art dense grid-based methods while achieving reduced memory consumption and improved efficiency.
+II Related Work
+Figure 2: GaussianFormer3D Overview.
+We first voxelize LiDAR point clouds to obtain non-empty voxel features for initializing the position and opacity of 3D Gaussians.
+Then LiDAR depth maps and camera feature maps are extracted respectively, and multiplied via outer product to construct a lifted 3D feature space.
+Gaussians are iteratively updated with sparse convolution, 3D deformable attention, and property refinement.
+Gaussians are eventually processed by a Gaussian-to-voxel splatting module to generate semantic occupancy.
 Multi-Modal Semantic Occupancy Prediction.
-Multi-modal occupancy methods generally outperform single-modal ones since different modalities can complement each other.
-Common multi-modal sensor configurations for semantic occupancy include LiDAR-camera [52, 66, 32, 3, 63, 60, 87], camera-radar [72, 45] and LiDAR-camera-radar [51].
-Among them, LiDAR-camera is the top-performing one as it combines LiDAR’s accurate depth and geometry sensing with camera’s powerful semantic perception.
-Similar to single-modal pipelines, existing LiDAR-camera occupancy networks mainly build on voxel-based [52, 66, 32, 87, 68] or BEV-based representations [60].
+Multi-modal occupancy prediction methods generally outperform single-modal approaches, as different modalities provide complementary information.
+Among them, LiDAR-camera fusion is the top-performing configuration, combining LiDAR’s accurate depth and geometry sensing with the powerful semantic understanding capability of cameras.
+Similar to single-modal pipelines, most LiDAR-camera occupancy prediction networks are also voxel-based [26, 34, 16, 45, 36].
 CONet [36] proposes a coarse-to-fine pipeline to sample 3D voxel features for refining the coarse occupancy prediction.
 Co-Occ [26] obtains multi-modal voxel features through a geometric and semantic-aware fusion module, and employs a NeRF-based implicit volume rendering regularization [25] to enhance the fused representation.
-OccGen [66] and OccMamba [32] encode multi-modal inputs to produce voxel fusion features, and then decode the features using diffusion denoising [17] and hierarchical Mamba modules [14] respectively to predict semantic occupancy.
-OccFusion [87] transforms LiDAR and camera inputs into multi-modal voxel features via 2D deformable attention [94] followed by an occupancy head.
+OccGen [34] and OccMamba [16] encode multi-modal inputs to produce voxel fusion features, and then decode the features using diffusion denoising and hierarchical Mamba modules, respectively.
+OccFusion [45] transforms LiDAR and camera inputs into multi-modal voxel features via 2D deformable attention [52].
 3D Gaussians for Autonomous Driving.
-Due to the inherent advantages of modeling scenes explicitly and continuously, 3D Gaussians [28] have been adopted as the scene representation over the traditional grid-based solutions in 3D semantic occupancy prediction [25, 23, 13] and 4D semantic occupancy forecasting [96].
-3D Gaussians also demonstrated their superiority in real-time image rendering and novel view synthesis and thus have been adopted for driving scene reconstruction and simulation [92, 56, 33, 78, 16].
-Furthermore, end-to-end autonomous driving [90] and visual pre-training [74, 85] utilize 3D Gaussians as the driving world representation for various downstream perception and planning tasks.
+Due to the inherent advantages of modeling scenes explicitly and continuously, 3D Gaussians [13] have been adopted as the scene representation over the traditional grid-based solutions in 3D semantic occupancy prediction [11, 9, 53].
+3D Gaussians also demonstrated their superiority in real-time image rendering and novel view synthesis, and thus have been adopted for driving scene reconstruction and simulation [50, 42, 8].
+Furthermore, end-to-end autonomous driving [48] and visual pre-training [41] utilize 3D Gaussians as the driving world representation for various downstream perception and planning tasks.
 However, these approaches are mainly designed for camera-only autonomous driving, neglecting the potential of multi-modal data in Gaussian initialization and updating.
 GSPR [29] proposes a Gaussian-based multi-modal place recognition algorithm, and SplatAD [8] designs the first 3D Gaussian splatting pipeline to render both LiDAR and camera data.
-In this work, we explore utilizing multi-modal data, especially from LiDAR and camera, to learn a better 3D Gaussian representation for more accurate and efficient semantic occupancy prediction.
-3 Method
+In this work, we explore utilizing multi-modal data, especially from LiDAR and camera sensors, to learn a fine-grained 3D Gaussian representation for more accurate and efficient semantic occupancy prediction.
+III Method
 The overview of GaussianFormer3D is presented in Fig. 2.
-3.1 Scene as 3D Gaussian Representation
-Semantic occupancy prediction aims to understand both semantic information and geometric structure of the environment.
-In the multi-modal scenario, given multi-view camera images =$\mathcal{I}=\{\mathbf{I}_{i}\}_{i=1}^{N_{c}}$ and LiDAR point cloud $\mathcal{P}=\{\mathbf{P}_{i}\}_{i=1}^{N_{p}}$, $\mathbf{P}_{i}=(x_{i},y_{i},z_{i},\eta_{i})$ containing the 3D position and intensity of the point, the goal is to predict the semantic occupancy grid $\mathbf{O}\in\mathcal{C}^{X\times Y\times Z}$, where $N_{c}$, $N_{p}$, $\mathcal{C}$ and $X\times Y\times Z$ represent the number of camera views, the number of LiDAR points, the set of semantic classes and the size of the voxel grid, respectively.
-Unlike uniform grids in traditional dense grid-based representations, the 3D Gaussian representation can adaptively depict the regions of interest due to the universal approximation capability of Gaussian mixtures [25, 11].
-Specifically, a scene is modeled as a set of 3D Gaussians $\mathcal{G}=\{\mathbf{G}_{i}\}_{i=1}^{N_{g}}$, where each Gaussian $\mathbf{G}_{i}$ is parameterized by its mean $\mathbf{m}_{i}\in\mathbb{R}^{3}$, rotation $\mathbf{r}_{i}\in\mathbb{R}^{4}$, scale $\mathbf{s}_{i}\in\mathbb{R}^{3}$, opacity $\sigma_{i}\in[0,1]$ and semantic label $\mathbf{c}_{i}\in\mathbb{R}^{|\mathcal{C}|}$.
-$N_{g}$ is the total number of Gaussians representing the scene.
+III-A Scene as 3D Gaussian Representation
+Semantic occupancy prediction aims to jointly predict the semantic information and geometric structure of the scene.
+Given multi-view images $\mathcal{I}=\{\mathbf{I}_{i}\}_{i=1}^{N_{c}}$ and LiDAR point cloud $\mathcal{P}=\{\mathbf{P}_{i}\}_{i=1}^{N_{p}}$, $\mathbf{P}_{i}=(x_{i},y_{i},z_{i},\eta_{i})$ containing the 3D location and intensity of each point, the goal is to predict the semantic occupancy grid $\mathbf{O}\in\mathcal{C}^{X\times Y\times Z}$, where $N_{c}$, $N_{p}$, $\mathcal{C}$ denote the number of camera views, the number of LiDAR points, and the set of semantic classes, and $X\times Y\times Z$ is the size of the voxel grid to be predicted.
+Unlike uniform grids in traditional grid-based representations, 3D Gaussians can adaptively represent the regions of interest due to the universal approximation capability of Gaussian mixtures [11].
+Specifically, a scene is modeled as a set of 3D Gaussians $\mathcal{G}=\{\mathbf{G}_{i}\}_{i=1}^{N_{g}}$, where $N_{g}$ is the total number of Gaussians in a scene.
+Each Gaussian $\mathbf{G}_{i}$ is parameterized by its mean $\mathbf{m}_{i}\in\mathbb{R}^{3}$, rotation $\mathbf{r}_{i}\in\mathbb{R}^{4}$, scale $\mathbf{s}_{i}\in\mathbb{R}^{3}$, opacity $\sigma_{i}\in[0,1]$ and semantic label $\mathbf{c}_{i}\in\mathbb{R}^{|\mathcal{C}|}$.
 The value of Gaussian $\mathbf{G}$ evaluated at location $\mathbf{x}$ can be calculated as:
@@
 $\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{\rm T}\mathbf{R}^{\rm T},\quad\mathbf{S}={\rm{diag}}(\mathbf{s}),\quad\mathbf{R}={\rm{q2r}}(\mathbf{r}),$ (2)
-[…]
+where $\mathbf{\Sigma}$, $\mathbf{R}$ and $\mathbf{S}$ denote the covariance matrix, rotation matrix and scale matrix.
+$\rm{diag}(\cdot)$ is the diagonal matrix construction and $\rm{q2r}(\cdot)$ is the quaternion-to-rotation transformation.
+By summing the contributions of all Gaussians at location $\mathbf{x}$, the occupancy prediction can be formulated as:
 $\hat{\mathbf{o}}(\mathbf{x};\mathcal{G})=\sum_{i=1}^{N_{g}}\mathbf{g}_{i}(\mathbf{x};\mathbf{m}_{i},\mathbf{s}_{i},\mathbf{r}_{i},\sigma_{i},\mathbf{c}_{i}).$ (3)
-Gaussian-to-voxel splatting module is designed to only aggregate Gaussians within a neighborhood of the targeted voxel instead of querying all Gaussians in a scene to improve efficiency and reduce unnecessary computation and storage [25].
+The Gaussian-to-voxel splatting module is designed to only aggregate Gaussians within the neighborhood of a targeted voxel instead of querying all Gaussians in a scene to improve efficiency and reduce unnecessary computation and storage [11].
 Thus, Eq.
@@
 During training, the Gaussian-based occupancy model is trained in an end-to-end manner, supervised by the ground truth semantic occupancy label $\mathbf{\bar{O}}\in\mathcal{C}^{X\times Y\times Z}$.
-Both cross entropy loss $L_{ce}$ and the lovasz-softmax loss $L_{lov}$ [2] are used for supervision.
-3.2 Voxel-to-Gaussian Initialization
-Two sets of 3D Gaussian features are adopted in our model following GaussianFormer [25].
+Both cross entropy loss $L_{ce}$ and the lovasz-softmax loss $L_{lov}$ are used for optimization.
+III-B Voxel-to-Gaussian Initialization
+Two sets of 3D Gaussian features are adopted following GaussianFormer [11].
 The first set consists of learnable Gaussian physical properties $\mathcal{G}=\{\mathbf{G}_{i}\in\mathbb{R}^{d}\}_{i=1}^{N_{g}}$ introduced in Sec.
-3.1, where $d=11+|\mathcal{C}|$, which are also our learning targets.
-The second set is non-learnable high-dimensional Gaussian features $\mathcal{Q}=\{\mathbf{Q}_{i}\in\mathbb{R}^{m}\}_{i=1}^{N_{g}}$, where $m$ is the feature dimension, serving as queries for the attention mechanism [65, 7, 94] and implicitly encode the spatial and semantic information during the Gaussian update.
-Previous work [25] randomly initializes the Gaussian physical properties during training, and optimizes these properties iteratively through multiple repetitive refinement modules.
-This design constrains Gaussians to learn complex 3D geometry information solely from 2D images, which will inevitably encounter inaccurate spatial modeling.
+III-A, where $d=11+|\mathcal{C}|$, which are also our learning targets.
+The second set is non-learnable high-dimensional Gaussian features $\mathcal{Q}=\{\mathbf{Q}_{i}\in\mathbb{R}^{m}\}_{i=1}^{N_{g}}$, where $m$ is the feature dimension, serving as queries for the attention mechanism [33] and implicitly encoding the spatial and semantic information during the Gaussian update.
+Previous work [11] randomly initializes the Gaussian physical properties, and optimizes these properties iteratively through multiple refinement modules.
+This design constrains Gaussians to learn complex 3D geometry information solely from 2D images, which inevitably encounters inaccurate spatial modeling.
 To resolve this issue, we propose a LiDAR-based voxel-to-Gaussian initialization strategy to initialize the mean and opacity of Gaussians with geometry priors from LiDAR data, as indicated in the dashed blue box in Fig.
@@
 Specifically, we first aggregate the most recent $N_{f}$ LiDAR scans into a combined point cloud $\mathcal{\bar{P}}=\{\mathcal{P}_{i}\}_{i=1}^{N_{f}}$.
-We voxelize the combined point cloud and compute the feature of each non-empty voxel as the mean position and intensity of all points within it [93].
+Then we voxelize the combined point cloud and compute the feature of each non-empty voxel as the mean position and intensity of all points within it.
 These LiDAR-based voxel features are then used to initialize the position and opacity of 3D Gaussians:
@@
 where $i\in\{1,...,N_{g}\}$ denotes the index of Gaussians to be initialized and $v\in\{1,...,N_{v}\}$ denotes the index of all non-empty voxels; $\mathcal{P}_{v}$ is the set of LiDAR points in $\mathcal{\bar{P}}$ falling into voxel $v$.
-When $N_{g}<N_{v}$, we randomly choose a subset of non-empty voxels with size of $N_{g}$ to initialize Gaussians, otherwise, we randomly select a subset of Gaussians with size of $N_{v}$ to be initialized with non-empty voxels.
-After initialization, we apply a 3D sparse convolution module to 3D Gaussians for self-encoding.
-The information and interactions of these Gaussians are efficiently extracted and aggregated through the sparse convolution operation for updating the Gaussian queries.
-3.3 LiDAR-Guided 3D Deformable Attention
-Lift, Splat, Shoot (LSS) [54] and 2D attention-based methods [94, 7] are widely adopted for feature lifting which transform multi-view 2D images into a 3D space to obtain lifted features.
+When $N_{g}<N_{v}$, we randomly choose a subset of $N_{g}$ non-empty voxels to initialize Gaussians, otherwise, a subset of $N_{v}$ Gaussians are randomly selected and initialized with non-empty voxels.
+After initialization, we apply a 3D sparse convolution module to the initialized 3D Gaussians for self-encoding.
+The features and interactions of Gaussians are efficiently extracted and aggregated through the sparse convolution for updating the Gaussian queries.
+III-C LiDAR-Guided 3D Deformable Attention
+Lift, Splat, Shoot (LSS) [28] and 2D attention-based methods [52] are widely adopted for feature lifting which transform multi-view 2D images into a 3D space to obtain lifted features.
 However, LSS suffers from excessive computational costs, hindering its application to multi-scale feature maps that are important for recognizing objects of various sizes.
-GaussianFormer [25] utilizes a 2D deformable attention to extract visual information from 2D images.
-Despite its efficiency, the 2D deformable attention-based method struggles with depth ambiguity.
-As multiple 3D reference points of different Gaussians can be projected to the same 2D position with similar sampling points in the 2D view, leading to the aggregation of the same 2D features for different 3D Gaussian queries.
-The underlying reason for this is the lack of accurate depth information during the feature lifting and aggregating process.
+GaussianFormer [11] utilizes a 2D deformable attention (Fig.
+3(a)) to extract visual information from 2D images.
+Despite its efficiency, it suffers from the depth ambiguity problem.
+As multiple 3D reference points from different Gaussians can be projected to the same 2D position with similar sampling points in the 2D view, this leads to ineffective aggregation, i.e., aggregating the same 2D features for different 3D Gaussian queries.
+The underlying reason for this is the lack of accurate depth information during the feature lifting and aggregating.
 A 3D deformable attention operator, namely DFA3D [17], is designed to mitigate the depth ambiguity problem by first expanding 2D feature maps into 3D using estimated depth [6] and then applying an attention mechanism [33] to aggregate features from the expanded 3D feature maps.
-However, the operator was originally designed for BEV-based 3D object detection, and relies on DepthNet [12] to estimate monocular depth.
+However, the operator is originally designed for BEV-based 3D object detection (Fig.
+3(b)), and relies on DepthNet [6] to estimate monocular depth.
 Inspired by DFA3D [17], we propose a LiDAR-guided 3D deformable attention mechanism for Gaussian-based semantic occupancy prediction, as illustrated in the dashed orange box in Fig.
@@
 We first form a unified LiDAR-camera 3D feature space $\mathbf{F}^{{\rm 3D}}$ by conducting outer product between the multi-scale depth maps $\mathbf{F}^{\rm d}$, generated from the LiDAR point cloud, and the multi-scale camera feature maps $\mathbf{F}^{\rm c}:\mathbf{F}^{\rm 3D}=\mathbf{F}^{\rm d}\otimes\mathbf{F}^{\rm c}$.
-For feature sampling, we design a two-stage key point sampling method to aggregate sufficient informative features for updating Gaussian queries.
+For feature sampling, we design a two-stage key point sampling method (Fig.
+3(c)) to aggregate sufficient informative features for updating Gaussian queries.
 First, we sample a group of 3D reference points $\mathcal{R}_{G}=\{\mathbf{m}_{i}=\mathbf{m}+\Delta\mathbf{m}_{i}|i=1,...,N_{R_{1}}\}$ for each Gaussian $\mathbf{G}$ by shifting its mean $\mathbf{m}$ with learned offsets $\Delta\mathbf{m}$.
 Then we project these 3D reference points into the fusion feature space $\mathbf{F}^{\rm 3D}$ with extrinsics $\mathcal{T}$ and intrinsics $\mathcal{K}$, where each projected reference point is positioned at $\mathbf{\bar{m}}_{i}=(u_{i},v_{i},d_{i})$.
+Figure 3: Comparison of different feature sampling methods.
 After projection, we further generate learnable sampling offsets $\Delta\mathbf{\bar{m}}_{ij}=(\Delta u_{ij},\Delta v_{ij},\Delta d_{ij})$ for each projected reference point $\mathbf{\bar{m}}_{i}$.
@@
 […]
-where $N_{R_{1}}$ and $N_{R_{2}}$ denote the number of sampling reference points for each Gaussian and for each projected 3D reference point, respectively.
-Finally, we update the Gaussian query $\mathbf{Q}$ with the weighted sum of aggregated LiDAR-camera fusion features $\Delta\mathbf{Q}$:
+where $N_{R_{1}}$ and $N_{R_{2}}$ denote the number of sampling points for each Gaussian and for each projected 3D reference point.
+Finally, the Gaussian query $\mathbf{Q}$ is updated with the weighted sum of aggregated LiDAR-camera fusion features $\Delta\mathbf{Q}$:
 $\Delta\mathbf{Q}=\frac{1}{N_{c}}\sum_{c=1}^{N_{c}}\sum_{i=1}^{N_{R_{1}}}\sum_{j=1}^{N_{R_{2}}}{\rm DFA}(\mathbf{Q},\boldsymbol{\pi}_{c}(\mathbf{\bar{m}}_{ij};\mathcal{T},\mathcal{K}),\mathbf{F}_{c}^{\rm 3D}),$ (6)
-where $\rm DFA(\cdot)$ and $\boldsymbol{\pi}_{c}(\cdot)$ represent the 3D deformable attention operation and the transformation from the Gaussian coordinate frame to $\mathbf{F}_{c}^{{\rm 3D}}$ coordinate frame generated from camera view $c$ respectively.
-After acquiring sufficient 3D geometric and semantic information through sparse convolution and 3D deformable attention, the Gaussian query $\mathbf{Q}$ is passed to a multi-layer perceptron (MLP), and decoded to refine the Gaussian property $\mathbf{G}$.
-We iteratively optimize the Gaussian properties with $B$ blocks of sparse convolution, 3D deformable attention, and refinement modules.
+where $\rm DFA(\cdot)$ and $\boldsymbol{\pi}_{c}(\cdot)$ represent the 3D deformable attention and the transformation from the Gaussian frame to $\mathbf{F}_{c}^{{\rm 3D}}$ frame generated from camera view $c$, respectively.
+After acquiring sufficient geometric and semantic information through sparse convolution and 3D deformable attention, the Gaussian query $\mathbf{Q}$ is passed to a multi-layer perceptron, and decoded to refine the Gaussian property $\mathbf{G}$.
+We iteratively optimize the properties with 4 blocks of sparse convolution, 3D deformable attention, and refinement modules.
+TABLE I: 3D semantic occupancy prediction results on nuScenes-SurroundOcc [38] validation set.
 Method Modality IoU $\uparrow$ mIoU $\uparrow$
@@
 Ours L+C 43.3 27.1 26.9 15.8 32.7 36.1 18.6 21.7 24.1 13.0 21.3 29.0 40.6 23.7 27.3 28.2 32.6 42.3
-Table 1: 3D semantic occupancy prediction results on SurroundOcc [70] validation set.
-The best is bolded and the second best is underlined.
-Our method achieves comparable overall performance with the state-of-the-art LiDAR-camera fusion-based methods, while surpasses them on classes of small objects, dynamic vehicles, and large surfaces.
+TABLE II: 3D semantic occupancy prediction results on nuScenes-Occ3D [32] validation set.
+* denotes training with camera visibility mask.
+(xf) denotes the number of history image frames used for temporal fusion.
 Method Modality mIoU $\uparrow$
@@
 MonoScene [3] C 6.1 1.8 7.2 4.3 4.9 9.4 5.7 4.0 3.0 5.9 4.5 7.2 14.9 6.3 7.9 7.4 1.0 7.7
-BEVDet [21] C 11.7 2.1 15.3 0.0 4.2 13.0 1.4 0.0 0.4 0.13 6.6 6.7 52.7 19.0 26.5 21.8 14.5 15.3
 BEVFormer [19] C 23.7 5.0 38.8 10.0 34.4 41.1 13.2 16.5 18.2 17.8 18.7 27.7 49.0 27.7 29.1 25.4 15.4 14.5
@@
 GaussianFormer* [11] C 35.5 8.8 40.9 23.3 42.9 49.7 19.2 24.8 24.4 22.5 29.4 35.3 79.0 36.9 46.6 48.2 38.8 33.1
-BEVDet4D* (2f) [20] C 39.3 9.3 47.1 19.2 41.5 52.2 27.2 21.2 23.3 21.6 35.8 38.9 82.5 40.4 53.8 57.7 49.9 45.8
 COTR* (2f) [22] C 44.5 13.3 52.1 32.0 46.0 55.6 32.6 32.8 30.4 34.1 37.7 41.8 84.5 46.2 57.6 60.7 52.0 46.3
@@
 Ours* L+C 46.4 9.8 50.0 31.3 54.0 59.4 28.1 36.2 46.2 26.7 40.2 49.7 79.1 37.3 49.0 55.0 69.1 67.6
-Table 2: 3D semantic occupancy prediction results on Occ3D [64] validation set.
-* denotes training with camera visibility mask.
-(xf) denotes the number of history image frames used for temporal fusion.
-The best is bolded and the second best is underlined.
-4 Experiments
-4.1 Datasets
+IV Experiments
+IV-A Datasets
 […]
-RELLIS-3D [27] dataset is a multi-modal off-road driving dataset containing RGB images, LiDAR point clouds, stereo images, GPS and IMU data.
+RELLIS-3D [12] dataset is a multi-modal off-road driving dataset collected by a Clearpath Warthog robot containing RGB images, LiDAR point clouds, stereo images, GPS and IMU data.
 WildOcc [44] provides the first off-road occupancy annotation on the RELLIS-3D, which are split into 7399/1249/1399 frames for training, validation and testing respectively.
-The annotation is in the range of $[-20\rm{m},0\rm{m}]\times[-10\rm{m},10\rm{m}]\times[-2\rm{m},6\rm{m}]$, where each voxel has a resolution of $0.2\mathrm{m}$ and labeled as one of 9 classes (7 semantics, 1 other class and 1 empty).
-WildOcc [83] is used to evaluate the performance of our model on off-road scenes with a LiDAR-monocular setting.
-4.2 Implementation and Evaluation Details
+The annotation is in the range of $20\rm{m},0\rm{m}]\times[-10\rm{m},10\rm{m}]\times[-2\rm{m},6\rm{m}]$, where each voxel has a resolution of $0.2\mathrm{m}$ and labeled as one of 9 classes (7 semantics, 1 other class and 1 empty).
+WildOcc [44] is used to evaluate the performance of our model in complex off-road environments and with a monocular-LiDAR sensor configuration.
+IV-B Implementation and Evaluation Details
 For camera branch, we set the resolution of input images as $900\times 1600$ for nuScenes [2] and $1200\times 1920$ for RELLIS-3D [12].
@@
 We employ these Gaussians to only model the occupied space, and leave the empty space to one fixed large Gaussian to improve efficiency [9].
-We train our model with an AdamW [42] optimizer with a weight decay of 0.01.
+We train our model with an AdamW optimizer with a weight decay of 0.01.
 The learning rates are set to $1\times 10^{-4}$ for nuScenes and $3\times 10^{-4}$ for RELLIS-3D, and decay with a cosine annealing schedule.
@@
 We use Intersection-over-Union (IoU) and mean Intersection-over-Union (mIoU) for evaluation metrics following MonoScene [3].
-See supplementary material for the calculation details of the metrics.
+TABLE III: 3D semantic occupancy prediction results on RELLIS3D-WildOcc [44] dataset.
 Method Mod. IoU $\uparrow$ mIoU $\uparrow$
@@
 Ours L+C 29.5 13.1 19.1 38.5 10.6 0.1 4.6 4.2 14.5
-Table 3: 3D semantic occupancy prediction results on WildOcc [83] validation and test sets.
-We show the percentages of classes in the ground truth occupied voxels for validation and test sets respectively.
-Results except for GaussianFormer [25] and ours are reported in WildOcc [83].
-(xf) denotes the number of history image frames used for temporal fusion.
-The best is bolded.
+TABLE IV: Efficiency comparison and ablation on number of Gaussians. Tested on one A40 GPU with one batch during inference.
 Method Mod. Query Form Query Number Lat. (ms) $\downarrow$ Mem. (GB) $\downarrow$ IoU $\uparrow$ mIoU $\uparrow$
-BEVFormer [37] C 2D BEV 200$\times$200 310 4.5 30.5 16.8
-TPVFormer [22] C 3D TPV 200$\times$(200+16+16) 320 5.1 30.9 17.1
-SurroundOcc [70] C 3D Voxel 200$\times$200$\times$16 340 5.9 31.5 20.3
+BEVFormer [19] C 2D BEV $\times200$ 310 4.5 30.5 16.8
+[…]
+SurroundOcc [38] C 3D Voxel $\times200$$\times$16 340 5.9 31.5 20.3
 25600 227 4.7 28.7 16.0
@@
 GaussianFormer-2 [9] C 3D Gaussian 25600 357 3.0 31.0 20.3
-50$\times$50$\times$4 532 7.6 33.3 21.2
-M-CONet [68] L+C 3D Voxel 100$\times$100$\times$8 670 7.8 39.2 24.7
-Co-Occ [52] L+C 3D Voxel 100$\times$100$\times$8 580 11.8 41.1 27.1
+$\times50$$\times$4 532 7.6 33.3 21.2
+M-CONet [36] L+C 3D Voxel $\times100$$\times$8 670 7.8 39.2 24.7
+Co-Occ [26] L+C 3D Voxel $\times100$$\times$8 580 11.8 41.1 27.1
 6400 415 4.9 39.6 21.4
@@
 Ours L+C 3D Gaussian 25600 555 5.5 43.3 27.1
-Table 4: Efficiency comparison of different methods and effect of Gaussian number on SurroundOcc [70] validation set.
-The results of our method are tested on one A40 GPU with batch size one during inference.
-The best is bolded within each modality.
-4.3 Quantitative Results
+IV-C Quantitative Results
 3D semantic occupancy prediction performance.
 We report the performance of GaussianFormer3D on SurroundOcc [38], Occ3D [32] and WildOcc [44] in Tab.
-1, Tab.
-2 and Tab.
-3, respectively.
+I, Tab.
+II and Tab.
+III, respectively.
 For on-road scenarios in Tab.
-1 and Tab.
-2, our method surpasses GaussianFormer [25] extensively on all classes, leading to overall 13.5 and 8.0 increases on the IoU and mIoU respectively on SurroundOcc [70] and 10.9 increase on the mIoU on Occ3D [64].
-Compared to state-of-the-art LiDAR-camera approaches [52, 68, 87], GaussianFormer3D achieves comparable overall performance while showing superior performance in predicting small objects (e.g., motorcycle, pedestrian), dynamic vehicles (e.g., car, construction vehicle, truck) and surrounding surfaces (e.g., manmade, vegetation) which are crucial classes for autonomous driving tasks.
+I and Tab.
+II, our method surpasses GaussianFormer [11] extensively on all classes, leading to overall 13.5 and 8.0 increases on the IoU and mIoU respectively on SurroundOcc [38] and 10.9 increase on the mIoU on Occ3D [32].
+Compared to state-of-the-art LiDAR-camera approaches [26, 36, 45], ours achieves best overall performance while showing superior performance in predicting small objects (e.g., motorcycle, pedestrian), dynamic vehicles (e.g., car, construction vehicle, truck) and surrounding surfaces (e.g., manmade, vegetation) which are crucial classes for autonomous driving tasks.
 This improvement is due to Gaussians’ universal approximating ability to model objects with flexible scales and shapes.
 For off-road results in Tab.
-3, our method with single-frame image input surpasses M-OFFOcc [83] using 4 sequential images by 1.1 in IoU and performs on par in mIoU.
+III, our method with single-frame image input surpasses M-OFFOcc [44] using 4 sequential images by 1.1 in IoU and performs on par in mIoU.
 Moreover, our method outperforms GaussianFormer [11] by 14.4 in IoU and 6.8 in mIoU on the test set, highlighting LiDAR’s role in understanding the geometry of complex off-road terrains.
-GaussianFormer3D excels in predicting regions with large surfaces, such as grass, tree, and puddle, while remaining suboptimal for subtle terrain variations like mud.
+Our method excels in predicting regions with large surfaces, such as grass, tree, and puddle, while remaining suboptimal for subtle terrain variations like mud.
 For barrier and rubble, their low occurrence in the test set (0.001% of occupied voxels) poses a challenge due to the lack of sufficient features for reliable prediction.
-See supplementary material for more analysis.
-Evaluation of model efficiency and effect of Gaussian number.
-We evaluate and compare the latency and memory consumption of our approach with other methods in Tab.
-4.
-Our model achieves multi-modal fusion-based prediction performance while maintaining approximately the same low memory usage as camera-only methods.
-Compared to Co-Occ [52], our method saves about 50% memory consumption, making it more suitable for running onboard on autonomous vehicles.
-Additionally, our approach employs only 25,600 Gaussians with 28 channels while Co-Occ [52] requires 80,000 queries with 128 channels to achieve similar performance demonstrating the potential of our method to enable more efficient communication for connected vehicles or multi-robot collaborations.
-The latency of our method is higher than that of vision-based pipelines, which is mainly due to the computation overhead introduced by 3D deformable attention operations.
+We further evaluate the model performance under different weather conditions in Tab.
+V.
+Ours shows a significant performance improvement over the baseline under extreme climate (rainy) and low lighting condition (night).
+Evaluation of model efficiency.
+We evaluate and compare the latency and average memory consumption of ours with other methods during testing in Tab.
+IV.
+Ours achieves multi-modal fusion-based prediction performance while maintaining approximately the same low memory usage as camera-only methods.
+Compared to Co-Occ [26], ours saves about 50% average memory consumption, making it more suitable for running onboard on autonomous vehicles.
+In addition, our approach employs only 25,600 Gaussians with 28 channels while Co-Occ [26] requires 80,000 queries with 128 channels to achieve similar performance, demonstrating the potential of our method to enable more efficient communication for connected vehicles or multi-robot collaborations.
+The latency of our method is higher than that of camera-only pipelines, which is mainly due to the computation overhead introduced by 3D deformable attention.
+Some LiDAR-camera methods [45, 44] are not compared due to lack of open-source code.
 We also examine the effect of the number of Gaussians on the model performance in Tab.
-4.
+IV.
 As the number of Gaussians increases, both latency and memory consumption rise, while the IoU and mIoU metrics are steadily improved.
-4.4 Ablation Study
-To break down the performance improvement brought by two designed modules, we conduct extensive ablation experiments to validate our design choices.
-The main ablation study is conducted in Tab.
-5.
-We observe that both the proposed voxel-to-Gaussian initialization and the LiDAR-guided 3D deformable attention modules contribute to the superior performance of GaussianFormer3D.
-The voxel-to-Gaussian initialization significantly improves the model’s ability to detect both small objects (e.g., pedestrian, traffic cone) and large surfaces (e.g., manmade structures, vegetation).
-This validates the effectiveness of multi-sweep LiDAR scans in providing Gaussians with accurate geometric information of occupied space.
-We also notice that LiDAR-guided 3D deformable attention mechanism enhances the model’s prediction ability on dynamic vehicles (e.g., bicycle, bus, car, motorcycle, trailer, truck) and near-road surfaces (e.g., drivable surface, flatten area, sidewalk, terrain) where objects detected by LiDAR points are visible to surrounding cameras.
-In these regions, the LiDAR points and corresponding image pixels are associated in the lifted 3D feature space, enabling the model to retrieve aggregated fusion features of on-road and near-road objects.
-Figure 3: Qualitative results on the on-road SurroundOcc [70] validation set.
+Figure 4: Qualitative results on the on-road nuScenes-SurroundOcc [38] validation set.
 Our multi-modal Gaussian-based occupancy method can capture both semantics information and geometry structure of the surroundings.
-Best viewed on screen and in color.
+Color legend is given in Tab.
+I.
+TABLE V: Performance on nuScenes-SurroundOcc [38] validation set under different weather and lighting conditions.
+[…]
+Method Modality [Uncaptioned image] Sunny [Uncaptioned image] Rainy [Uncaptioned image] Day [Uncaptioned image] Night [Uncaptioned image] Sunny [Uncaptioned image] Rainy [Uncaptioned image] Day [Uncaptioned image] Night
+GaussianFormer [11] C 29.6 27.5 30.3 19.5 18.9 18.0 19.2 9.3
+GaussianFormer3D L+C 43.6 (+14.0) 41.6 (+14.1) 43.6 (+13.3) 40.5 (+21.0) 27.3 (+8.4) 25.2 (+7.2) 27.4 (+8.2) 15.5 (+6.2)
+TABLE VI: Ablation study of proposed modules evaluated on nuScenes-SurroundOcc [38] validation set.
+Voxel-to-Gaussian and LiDAR-Guided 3D Deformable Attention are abbreviated as V2G and DFA respectively.
 Model V2G DFA IoU $\uparrow$ mIoU $\uparrow$
@@
 GaussianFormer3D 43.3 27.1 26.9 15.8 32.7 36.1 18.6 21.7 24.1 13.0 21.3 29.0 40.6 23.7 27.3 28.2 32.6 42.3
-Table 5: Ablation study of proposed modules evaluated on SurroundOcc [70] validation set. Voxel-to-Gaussian and LiDAR-Guided 3D Deformable Attention are abbreviated as V2G and DFA respectively.
+TABLE VII: Ablation study of module design choices on the nuScenes-SurroundOcc [38] validation set.
+(a) Ablation study of Gaussian initialization strategies. PM-
+Point denotes probabilistic modeling with point cloud in [9].
 […]
@@
 V2G 40.7 25.8
-(a) Ablation study of initialization strategy for V2G. PM-Point denotes probabilistic modeling with point cloud reported in [23].
+(b) Ablation study of LiDAR voxel size for V2G. The unit of length is $\mathrm{m}$. We set the height of all the voxels as $0.2\mathrm{m}$.
 […]
@@
 V2G 40.7 25.8
-(b) Ablation study of LiDAR voxel size for V2G. The unit of length is $\mathrm{m}$. We set the height of all the voxels as $0.2\mathrm{m}$.
+(c) Ablation study of offset sampling methods for DFA. We run experiments with applying learnable offset sampling before and after projecting Gaussians into the lifted 3D feature space.
 […]
@@
 DFA 40.7 26.4
-(c) Ablation study of offset sampling methods for DFA. We run experiments with applying learnable offset sampling before and after projecting Gaussians into the lifted 3D feature space.
+(d) Ablation study of feature lifting and aggregating methods for DFA. We concatenate LiDAR sparse and dense depth maps with RGB features respectively to conduct 2D deformable attention.
 […]
@@
 DFA 40.7 26.4
-(d) Ablation study of feature lifting and aggregating methods for DFA.
-We concatenate LiDAR sparse and dense depth maps with RGB features respectively to conduct 2D deformable attention-based Gaussian update.
-Table 6: Ablation study of module design choices on the SurroundOcc [70] validation set. Voxel-to-Gaussian and LiDAR-Guided 3D Deformable Attention are abbreviated as V2G and DFA respectively.
-Figure 4: Qualitative results on the off-road WildOcc [83] test set.
-Our multi-modal Gaussian-based occupancy method can outperform the ground truth (as shown in the first row) and predict classes such as puddle that are vital for off-road autonomous driving (as shown in the second row).
-Best viewed on screen and in color.
-Figure 5: Visualization comparison with GaussianFormer [25] on SurroundOcc [70].
+Figure 5: Qualitative results on the off-road RELLIS3D-WildOcc [44] test set.
+Our method can outperform the GT (first row) at some regions and predict classes such as puddle that are vital for off-road autonomous driving (second row).
+Color legend is given in Tab.
+III.
+Figure 6: Visualization comparison with GaussianFormer [11] on nuScenes-SurroundOcc [38].
 By incorporating LiDAR, our method can obtain Gaussians with more adaptive scales and shapes, resulting in more accurate semantic predictions and delicate geometry details.
-Figure 6: Multi-resolution occupancy prediction using same 3D Gaussians.
-Leveraging the continuity of Gaussians, our method enables multi-resolution prediction, yielding more accurate and smoother occupancy at higher resolution in certain regions.
+Figure 7: Multi-resolution occupancy prediction of 3D Gaussians.
+IV-D Ablation Study
+We conduct extensive ablation experiments to validate our design choices.
+The main ablation study is conducted in Tab.
+VI.
+We observe that both the proposed voxel-to-Gaussian initialization and the LiDAR-guided 3D deformable attention modules contribute to the superior performance of our method.
+The voxel-to-Gaussian initialization significantly improves the model’s ability to detect both small objects (e.g., pedestrian, traffic cone) and large surfaces (e.g., manmade, vegetation).
+This validates the effectiveness of multi-sweep LiDAR scans in providing Gaussians with accurate geometric information of occupied space.
+We also notice that LiDAR-guided 3D deformable attention mechanism enhances the model’s prediction ability on dynamic vehicles (e.g., bicycle, bus, car, motorcycle, trailer, truck) and near-road surfaces (e.g., drivable surface, flatten area, sidewalk, terrain) where objects detected by LiDAR points are visible to surrounding cameras.
+In these regions, the LiDAR points and corresponding image pixels are associated in the lifted 3D feature space, enabling the model to retrieve aggregated fusion features of on-road and near-road objects.
 Voxel-to-Gaussian Initialization.
 We first compare different levels of LiDAR features used for initializing Gaussian properties in Tab.
-6(a).
+VII(a).
 The improvement achieved with the multi-sweep voxel feature is significantly greater than that of the single-sweep point feature and the point cloud probabilistic modeling strategy used in GaussianFormer-2 [9], which validates the effectiveness of our proposed module.
 We further conduct an ablation study on the size of LiDAR voxel in initialization in Tab.
-6(b).
+VII(b).
 As the voxel size decreases, the model performance slightly improves.
@@
 We first study the effect of the two-stage offset sampling strategy in Tab.
-6(c).
-We observe that applying learnable offset sampling both before and after projection achieves higher performance than single-stage sampling, which validates that our two-stage sampling method can aggregate sufficient informative features for refining Gaussians.
-We also compare different feature aggregating methods for deformable attention in Tab.
-6(d), including 3D deformable attention, 2D deformable attention with concatenated LiDAR sparse depth map and with completed dense depth map [29].
-The results validate our design choice of 3D deformable attention.
-4.5 Qualitative Results
-We visualize 3D Gaussians and occupancy to qualitatively verify the effectiveness of GaussianFormer3D for on-road scenes in Fig.
-3.
+VII(c).
+We observe that applying learnable offset sampling both before and after projection achieves higher performance than single-stage sampling, which validates our two-stage sampling method can aggregate sufficient informative features for refining Gaussians.
+We also compare different feature aggregating methods in Tab.
+VII(d), including 3D deformable attention, 2D deformable attention with concatenated LiDAR sparse depth map and with completed dense depth map [14].
+The results validate our final choice.
+IV-E Qualitative Results
+We visualize 3D Gaussians and occupancy to qualitatively verify the effectiveness of our method for on-road scenes in Fig.
+4.
 Our method can accurately predict both semantics and fine-grained geometry of the surrounding environments.
-In some cases, it even outperforms the ground truth by correctly completing occupancy in regions that lack semantic annotations.
+In some cases, it even outperforms the GT by correctly completing occupancy in regions that lack semantic annotations.
 Qualitative results of our method on off-road scenes are given in Fig.
-4.
+5.
 Our method is able to predict semantic occupancy for classes like mud and puddle, which are essential for achieving safe and effective off-road autonomous driving.
 We further compare our approach with GaussianFormer [11] in Fig.
-5.
+6.
 The Gaussians in our method are more adaptive in scales and shapes, precisely appearing in the occupied regions of objects in both long-range and short-range areas, aided by the LiDAR sensor.
@@
 This property enables more accurate and smoother prediction for certain areas when inferred at a higher resolution, as demonstrated in Fig.
-6.
-Please see more qualitative results in the supplementary material.
-5 Conclusion
-In this paper, we proposed GaussianFormer3D, a novel multi-modal semantic occupancy prediction framework that builds on 3D Gaussian scene representation and 3D deformable attention.
+7.
+V Conclusion
+In this paper, we proposed GaussianFormer3D, a novel multi-modal semantic occupancy prediction framework that builds on 3D Gaussian scene representation.
 We introduced a voxel-to-Gaussian initialization strategy to endow 3D Gaussians with accurate geometry priors from LiDAR data.
-We also designed a LiDAR-guided 3D deformable attention mechanism to refine 3D Gaussians with LiDAR-camera fusion feature in a lifted 3D space.
-Extensive experiments show the effectiveness of GaussianFormer3D in achieving accurate and fine-grained semantic occupancy prediction.
-In the future, we plan to explore the application of our multi-modal 3D Gaussian scene representation for multi-robot coordination.
-Supplementary Material
-Appendix A Related Work
-Single-Modal Semantic Occupancy Prediction.
-Semantic occupancy prediction, also known as semantic scene completion, was firstly proposed to predict the occupancy and semantic states of all voxels within a pre-defined range for indoor scenarios [86, 34, 61, 9].
-The vision-based and LiDAR-based pipelines are first explored, and further extended to outdoor environments, especially the autonomous driving scenes [4, 1, 39, 62].
-The key to this challenging task is to learn an informative and compact scene representation, to which the common solutions mainly include 3D voxel-based representation [93, 77], 2D BEV-based representation [37, 79] and TPV-based (Tri-Perspective View) representation [22].
-Voxel-based methods utilize discretized voxels to represent a 3D space, and transform semantic and geometric information from 2D images [5, 26, 36, 43, 48, 69, 70, 88, 18, 91, 47, 84, 53] or 3D point clouds [58, 76, 57, 10, 80, 73, 46, 6, 41, 75, 50] to voxel feature vectors.
-These approaches can effectively capture complicated structures but incur high storage and computational costs due to the dense nature of voxel-based representations.
-Moreover, a significant portion of these costs is unnecessary given the abundance of empty voxels in autonomous driving scenes.
-BEV-based representation has also been applied to semantic occupancy prediction [38, 82, 24] after its success in 3D object detection [37, 79].
-TPV-based representation [22] extends BEV with two additional perpendicular planes and has been used in both vision-based [24, 22] and LiDAR-based [95] semantic occupancy networks.
-Although BEV-based and TPV-based methods mitigate the redundancy issue caused by the empty grids, their compression schemes inevitably lead to information loss.
-3D Gaussian splatting [28] provides an expressive and compact scene representation, which has been explored in vision-based 3D semantic occupancy prediction [25, 23, 13].
-GaussianFormer [25] represents an autonomous driving scene as a set of 3D Gaussians, designed to only model the object-occupied regions, which significantly improved prediction efficiency and reduced memory consumption.
-Appendix B WildOcc Dataset
-Figure 7: Distribution of categories across the WildOcc [83] train, validation, and test sets. The bar charts illustrate the percentage of each class within the respective dataset splits.
-Figure 8: Qualitative results of GaussianFormer3D on the WildOcc [83] validation and test sets.
-In the off-road scenes, our model can still capture the geometry structures and semantic information of large continuous surfaces (rubble, grass) or small irregular objects (barrier, mud), as shown in the red boxes.
-Best viewed on screen and in color.
-[…]
-Based on the split files from RELLIS-3D [27], we filter out samples in the WildOcc that lack corresponding semantic occupancy labels, LiDAR poses, or ground truth depth, obtaining 7399, 1249 and 1399 samples for training, validation and testing, respectively.
-The test set originally contains 11 classes in total.
-However, to ensure comparability with the baselines in [83], we follow the same class selection strategy, considering semantic classes including grass, tree, bush, barrier, puddle, mud, and rubble, while grouping the remaining labels into the others category.
-The percentage distribution of each class is presented in Fig.
-7.
-A key challenge of this dataset is the significant distribution shift between the training/validation sets and the test set.
-Notably, as shown in Fig.
-7, the barrier and rubble classes exhibit extremely low occurrence in the test set compared to other categories.
-Specifically, the test set contains only 273 and 233 voxels for barrier and rubble, respectively, accounting for merely 0.0008% and 0.0007% of the occupied voxels.
-This severe class imbalance poses challenges for model performance, as limited training data can hinder the accurate prediction of these underrepresented categories.
-Additionally, the insufficiency of captured features further complicates the prediction process, making it even more challenging to reliably infer these classes.
-Appendix C Evaluation Metrics
-For evaluation metrics, we utilize Intersection-over-Union (IoU) and mean Intersection-over-Union (mIoU) following MonoScene [5]. The IoU and mIoU can be computed as:
-$\mathrm{IoU}$ $=\frac{TP_{\neq c_{0}}}{TP_{\neq c_{0}}+FP_{\neq c_{0}}+FN_{\neq c_{0}}},$ (7)
-$\mathrm{mIoU}$ $=\frac{1}{|\mathcal{C}^{\prime}|}\sum_{i\in\mathcal{C}^{\prime}}{\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}},$ (8)
-where $TP$, $FP$, $FN$, $c_{0}$ and $\mathcal{C}^{\prime}$ represent the number of true positive, false positive, false negative predictions, the empty class and the set of non-empty classes respectively.
-Figure 9: Qualitative results of GaussianFormer3D on the Occ3D [64] validation set.
-Our method is able to recover the fine-grained geometry structures and accurate semantics of the on-road scenes, and even outperforms the ground truth by completing the occupancy at some areas without annotations.
-Best viewed on screen and in color.
-Figure 10: Multi-resolution occupancy predicted by the same 3D Gaussians on the SurroundOcc [70] validation set.
-The regions within the red boxes show that the high-resolution prediction generated by the same 3D Gaussians can achieve more accurate and smoother results, compared to the low-resolution one and the occupancy ground truth.
-Best viewed on screen and in color.
-Appendix D Implementation Details
-For multi-view images, we only apply photometric distortion augmentation, image normalization and image padding.
-During evaluation, we do not apply any test time augmentation technique.
-For the 3D deformable attention, DFA3D [31] has proved that the trilinear interpolation in the 3D feature space can be transformed into a depth-weighted bilinear interpolation, which means that in implementation, 3D deformable attention can be transformed into a depth-weighted 2D deformable attention to simultaneously maintain theoretical equivalence and improve efficiency.
-Hence we take advantage of the depth-weighted 2D deformable attention operator implemented in CUDA by DFA3D [31] to conduct 3D deformable attention.
-Appendix E Supplementary Experiments
-E.1 Qualitative Results
-Qualitative results on the WildOcc [83] validation and test sets.
-We present visualization results of our model on the WildOcc [83] validation and test sets in Fig.
-8.
-Two representative samples from both the validation and test sets are selected for illustration.
-Notably, all sequences in the validation and test sets are completely unseen during training.
-The class distribution has been reported in Appendix B.
-In the first validation sample (top row), our method successfully predicts the rubble region (highlighted in red boxes), which corresponds to the pile of rocks in the camera view.
-In the second validation sample, our model effectively captures the barrier structures in the scene.
-These results indicate that our model can learn semantic occupancy even with limited training data.
-In the test set samples (bottom two rows), our method accurately reconstructs large continuous surfaces, such as grass and trees, and effectively detects irregular objects and terrains, such as bush and mud, as shown in the red boxes.
-Overall, our predictions exhibit a high degree of similarity to the ground truth, especially in preserving the geometric structures of various terrain types, including grass, tree, rubble, and barrier.
-Qualitative results on Occ3D [64] validation set.
-We provide visualization results of our model on the Occ3D [64] validation set in Fig.
-9.
-Our method achieves fine-grained semantic occupancy prediction, and even occasionally outperforms the ground truth.
-Even though the occupancy ground truth is sparse and missing at some regions, our model is able to accurately recovering semantic information of these unlabeled areas.
-Figure 11: Visualization comparison of different Gaussian initialization strategies on the SurroundOcc [70] validation set.
-With LiDAR voxel features as priors, our 3D Gaussians precisely appear at the regions of interests, which paves the way for the following 3D deformable attention-based Gaussian update.
-[…]
-Method Modality Sunny Rainy Day Night Sunny Rainy Day Night
-GaussianFormer [25] C 29.6 27.5 30.3 19.5 18.9 18.0 19.2 9.3
-GaussianFormer3D L+C 43.6 (+14.0) 41.6 (+14.1) 43.6 (+13.3) 40.5 (+21.0) 27.3 (+8.4) 25.2 (+7.2) 27.4 (+8.2) 15.5 (+6.2)
-Table 7: 3D semantic occupancy prediction results on SurroundOcc [70] validation set for different weather and lighting conditions.
-We quantitatively show the improvement made by GaussianFormer3D over the baseline GaussianFormer.
-Multi-resolution semantic occupancy prediction on SurroundOcc validation set.
-We provide more qualitative results of multi-resolution semantic occupancy prediction in Fig.
-10.
-As shown in red boxes in Fig.
-10, we observe that the high-resolution prediction achieves more accurate and smoother occupancy at some regions.
-Although Gaussians are supervised with the fixed-resolution occupancy ground truth, multi-resolution predictions can still be generated by the same Gaussian representation due to its continuous modeling property.
-This advantage eradicates the need of training several models to predict occupancy of different resolutions, and keeps the high prediction accuracy with the same Gaussians, which significantly shortens the algorithm deployment time and saves the computation resources.
-Visualization comparison of different Gaussian initialization strategies.
-We provide qualitative results of visualization comparison of different Gaussian initialization strategies in Fig.
-11.
-Compared to random initialization, our designed voxel-to-Gaussian (V2G) initialization strategy is able to endow the 3D Gaussians with accurate position and geometry priors from the LiDAR data at the beginning of training, which paves the way for the following 3D deformable attention-based Gaussian refinement.
-E.2 Quantitative Results
-Model performance under different weather conditions.
-To break down the performance improvement under different weather conditions, we group the scenes in the nuScenes [4] dataset based on climate and lighting conditions.
-We demonstrate the results of both GaussianFormer3D and GaussianFormer [25] in Tab.
-7.
-Compared to the baseline GaussianFormer, our approach presents increased performance across all weather conditions, which validates the effectiveness and robustness of our model.
-Due to the introduction of LiDAR sensor, GaussianFormer3D shows a significant performance improvement over the camera-only baseline under extreme climate (rainy) and low lighting condition (night).
-Both experiments are conducted with 25,600 Gaussians for a fair comparison.
+We also designed a LiDAR-guided 3D deformable attention mechanism to refine 3D Gaussians with LiDAR-camera fusion feature.
+Extensive experiments show its effectiveness in achieving accurate and fine-grained semantic occupancy prediction.
+However, our model is limited to fully-supervised manner, which requires densely annotated occupancy labels for training.
+In the future, we will explore its self-supervised variant and its application for multi-robot coordination.